### v1

# Stability of the expanding region of Kerr de Sitter spacetimes

Abstract
We prove the nonlinear stability of the cosmological region of Kerr de Sitter spacetimes. More precisely, we show that solutions to the Einstein vacuum equations with positive cosmological constant arising from data on a cylinder that is uniformly close to the Kerr de Sitter geometry (with possibly different mass and angular momentum parameters at either end) are future geodesically complete and display asymptotically de Sitter-like degrees of freedom. The proof uses an ADM formulation of the Einstein equations in parabolic gauge. Together with a well-known theorem of Hintz-Vasy [Acta Math. 220 (2018)], our result yields a global stability result for Kerr de Sitter from Cauchy data on a spacelike hypersurface bridging two black hole exteriors.
1 Introduction
In the presence of a cosmological constant $\Lambda>0$, the Einstein vacuum equations take the form
${\bm{\mathrm{Ric}}}[{\bm{g}}]=\Lambda{\bm{g}}\,,$ (1.1)
where ${\bm{\mathrm{Ric}}}[{\bm{g}}]$ is the Ricci curvature of the spacetime metric ${\bm{g}}$ on a $3+1$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$. Since the sum of the sectional curvatures $\bm{K}_{i}$ is negative, [^1] Here $(e_{0},e_{1},e_{2},e_{3})$ is an orthonormal frame, $e_{0}$ is time-like, and $\bm{K}_{i}=\bm{R}(e_{0},e_{i},e_{0},e_{i})$ are the sectional curvatures associated to the planes spanned by $e_{0}$ and $e_{i}$.
${\bm{\mathrm{Ric}}}[{\bm{g}}](e_{0},e_{0})=\sum_{i=1}^{3}\,\bm{K}_{i}=\Lambda\>{\bm{g}}(e_{0},e_{0})<0$
solutions to (1.1) may exhibit expansion in all directions. The simplest example of a spacetime with this property is de Sitter space, which models a spatially closed expanding universe with topology $\mathbb{S}^{3}\times\mathbb{R}$. The exact solutions to (1.1) that motivate this paper are the Kerr de Sitter spacetimes $(\mathcal{M},{\bm{g}}_{\mathcal{K}_{a,m}})$. In addition to black hole interior and exterior regions, they contain a spatially open expanding, or cosmological region, with topology ×$\mathbb{R}\times\mathbb{S}^{2}\times\mathbb{R}$ [^2] Carter gives an excellent discussion of the maximal extension of Kerr de Sitter in [Car09]. For an introduction to the global geometry of the cosmological region, specifically in the context of the Cauchy problem see [Sch15, Sch22].
It is known since the work of Friedrich that de Sitter space is stable as a solution to (1.1): Small perturbations of the initial data on $\mathbb{S}^{3}$ lead to future geodesically complete spacetimes [Fri86]. The proof demonstrates in particular the existence of asymptotic functional degrees of freedom, but it does not apply to Kerr de Sitter spacetimes. [^3] In [Fri86] a conformal transformation is used to pass from (1.1) to the conformal field equations which turn out to be regular at the future boundary. In this way, Friedrich was able to reduce the global stability problem for de Sitter to a local in time problem, and identify the asymptotic degrees of freedom with the data on the conformal boundary. In Kerr de Sitter, the desired conformal transformation fails to be regular at $\iota^{+}$, and this approach is limited to spatially compact subsets of the cosmological region, and has been applied away from the endpoints in [MVK23, GVK17]. The exterior of slowly rotating Kerr de Sitter black holes have been proven to be asymptotically stable in a series of influential papers by Hintz and Vasy [Vas13, HV16, HV18]: In the domain bounded by the event and cosmological horizons, small perturbations settle down exponentially fast to a nearby member of the Kerr de Sitter family. In this paper, we complete the proof of the global nonlinear stability of Kerr de Sitter, proving that the cosmological region is stable:
Small perturbations on ×$\mathbb{R}\times\mathbb{S}^{2}$ which converge exponentially fast at both ends to nearby members of the Kerr de Sitter family lead to future geodesically complete solutions to (1.1) which display asymptotically de Sitter-like degrees of freedom.
We proceed with the precise statements.
Kerr de Sitter metric.
Recall the Penrose diagram of the Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ in Fig. 1, for small angular momentum $a^{2}\ll m^{2}\ll\Lambda^{-4}$. The cosmological region $\mathcal{R}$ lies to the future of the black hole exteriors $\mathcal{S}_{1}$, and $\mathcal{S}_{2}$, and is separated from these by the cosmological horizons $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. The conformal boundary at infinity is denoted by $\mathcal{I}^{+}$, and the black hole regions $\mathcal{B}_{1}$, and $\mathcal{B}_{2}$ are in the complement of its past.
Figure 1: Penrose diagram of Kerr de Sitter geometry.
The conformal diagram in Fig. 1 depicts various level sets of a function $r$, which are timelike, spacelike, or null depending on the values of the polynomial
$\Delta_{r}=(r^{2}+a^{2})\Bigr{(}1-\frac{\Lambda}{3}r^{2}\Bigr{)}-2mr\,,$ (1.2)
which may be positive, negative, or zero, respectively [Car09, GH77]. In the cosmological region $r$ is a time-function, and the metric takes the form
${\bm{g}}_{\mathcal{K}_{a,m}}=-\Phi_{\mathcal{K}_{a,m}}^{2}\mathrm{d}r^{2}+(g_{\mathcal{K}_{a,m}})_{r}\,,\qquad\Phi_{\mathcal{K}_{a,m}}^{-2}=\frac{\Lambda}{3}r^{2}-1+\mathcal{O}(r^{-1})\,,$ (1.3)
where $(g_{\mathcal{K}_{a,m}})_{r}$ is a Riemannian metric. Using a reparametrization of time
$r=e^{Hs}\,,\qquad H=\sqrt{\frac{\Lambda}{3}}\,,$ (1.4)
that leads to an expression for the metric that is more common in cosmology, [^4] In particular, in comparison to FLRW spacetimes; see for instance [RS13, Fou22] and references therein. we note that ${\bm{g}}_{\mathcal{K}_{a,m}}$ can then be expressed in coordinates so that in $\mathcal{R}$:
${\bm{g}}_{\mathcal{K}_{a,m}}=-\Phi_{\mathcal{K}_{a,m}}^{2}\mathrm{d}s^{2}+(g_{s,\mathcal{K}_{a,m}})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}\,,$ (1.5)
$\Phi_{\mathcal{K}_{a,m}}^{2}=1+\mathcal{O}(e^{-2Hs})\,,\qquad(g_{s,\mathcal{K}_{a,m}})_{ij}=\mathcal{O}(e^{2Hs})\,;$ (1.6)
[…]
Cauchy problem for Kerr de Sitter.
 [^6] For an introduction to the global stability problem for Kerr de Sitter see [HV18, Sch22]; also [DR13, Ch. 6].
Consider a spacelike hypersurface $\Sigma$ in Schwarzschild de Sitter as depicted in Fig. 2. The nonlinear stability result of Hintz and Vasy in [HV18] implies that for initial data $(g,k)$ close to the data induced by a Schwarzschild de Sitter metric ${\bm{g}}_{\mathcal{K}_{0,m}}$, the solution to (1.1) converges to nearby members of the Kerr de Sitter family in both $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$: [^7] The statement applies independently to $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ by domain of dependence. The result is obtained in a generalized harmonic gauge, which is itself determined dynamically together with the final states ${\bm{g}}_{\mathcal{K}_{a_{i},m_{i}}},i=1,2$. The specific gauge used in [HV18] is not relevant for the following, except that the time variable $t$ which is used to express the rate of convergence is comparable to Schwarzschild de Sitter time, $T(t)=1$ where $\mathcal{L}_{T}{\bm{g}}_{\mathcal{K}_{0,m}}=0$.
${\bm{g}}-{\bm{g}}_{\mathcal{K}_{a_{i},m_{i}}}=\mathcal{O}(e^{-\alpha t})\quad\text{: in }\mathcal{S}_{i}\,,\qquad i=1,2\,,$ (1.7)
for some parameters $(a_{1},m_{1})$ and $(a_{2},m_{2})$ with $\sum_{i=1}^{2}|a_{i}|+|m_{i}-m|\ll 1$, and $\alpha>0$. In fact, the stability result (1.7) holds on a domain that goes beyond the event and cosmological horizons, uniformly in $r$ [HV18, Theorem 1.1]. Therefore the existence of a development is known on a domain (whose future boundary is indicated by the dash dotted lines in Figure 2) which contains a level set $\Sigma_{s_{0}}\simeq\mathbb{R}\times\mathbb{S}^{2}$ of $r$, with the property that the geometric data on $\Sigma_{s_{0}}$ converges along one end to that induced by ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$, and that induced by ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$ along the other. This is the initial data for the evolution problem we consider.
Figure 2: Cauchy problem for Kerr de Sitter.
Geometric set-up for the main theorem.
We establish the existence of a spacetime $(\mathcal{R},{\bm{g}})$ which is foliated by the level sets of a time function $s$,
$\mathcal{R}=\bigcup_{s\in[s_{0},\infty)}\Sigma_{s}\,,\qquad\Sigma_{s}\simeq\mathbb{R}\times\mathbb{S}^{2}\,,$ (1.8)
with each leaf $\Sigma_{s}$ being diffeomorphic to a cylinder ×$\mathbb{R}\times\mathbb{S}^{2}$. We choose coordinates so that
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+(g_{s})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (1.9)
where $\Phi$ is the lapse function of the foliation, and $g_{s}$ is a Riemannian metric on $\Sigma_{s}$; see Section 2.1. The foliation is determined by a choice of the lapse function which we take to be the solution of a parabolic PDE; [^8] The PDE satisfied by the lapse function is the consequence of a geometric condition that involves a reference metric; see (1.12) below. We defer the derivation of the PDE to Section 2.2. The specific gauge choice is of course central to the global existence proof in this setting. For a broader discussion of the gauge in relation to other works in the literature see Remark 1.5 below. see Section 2.2.
Given the differentiable structure of $\mathcal{R}$, we can view ${\bm{g}}_{\mathcal{K}_{a,m}}$ as a family of metrics on $\mathcal{R}$. Moreover
$\widetilde{{\bm{g}}}=\chi(t){\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}+(1-\chi(t)){\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$ (1.10)
is a metric on $\mathcal{R}$, where $\chi$ is a smooth cutoff function on $\mathbb{R}$ with $\chi(t)=0$ for $t\leq-1$, and $\chi(t)=1$ for $t\geq 1$. With this choice a reference metric, which in the above coordinates again takes the form $\widetilde{{\bm{g}}}=-\widetilde{\Phi}^{2}\mathrm{d}s^{2}+(\widetilde{g}_{s})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ we can define
$\widehat{\Phi}=\Phi-\widetilde{\Phi}\,,\quad\widehat{g}=g-\widetilde{g}\,,\quad\>\widehat{k}=k-\widetilde{k}\,,$ (1.11)
where $g$, $\widetilde{g}$, and $k$, $\widetilde{k}$ are the first and second fundamental forms of ${\bm{g}}$, and $\widetilde{{\bm{g}}}$ on $\Sigma_{s}$, respectively.
Hyperbolicity and Energies.
In Section 2.2 we cast the Einstein equations (1.1) as a system of first order variation equations for $\widehat{g}$, and $\widehat{k}$, (and the renormalised Christoffel symbols $\widehat{\Gamma}=\Gamma-\widetilde{\Gamma}$). We show that in the gauge
$\Phi-\widetilde{\Phi}=\operatorname{tr}_{g}k-\operatorname{tr}_{\widetilde{g}}\widetilde{k}\,,$ (1.12)
the system of equations is essentially symmetric hyperbolic,  [^9] The hyperbolic structure of the system is seen only in conjunction with the constraint equations; see in particular Section 5.1. in the sense that we have an energy identity for the system derived in Section 5. The energies we use for the global existence argument are based on standard higher order Sobolev norms on $\Sigma_{s}$:
$\mathcal{E}_{N}(s)=\|\widehat{g}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{H^{N}(\Sigma_{s},g)}\big{)}$ (1.13)
While suppressed from the notation, the norms in the Sobolev spaces $H^{N}(\Sigma_{s},g)=H^{N}_{\alpha_{1},\alpha_{2}}(\Sigma_{s},g)$ are weighted to incorporate exponential decay towards the two ends of the cylinder; see Section 4.14.2. For fixed $\alpha_{1}\geq 0$, and $\alpha_{2}\geq 0$ in the definition of the norms, the Sobolev embedding reads; cf. Figure 3:
$\|(e^{\alpha_{1}t}+e^{-\alpha_{2}t})\mathcal{T}\|_{W^{N,\infty}(\Sigma_{s},g)}\leq C\|\mathcal{T}\|_{H_{\alpha_{1},\alpha_{2}}^{N+2}(\Sigma_{s},g)}$ (1.14)
Note that $\mathcal{E}_{N}(s)$ in (1.13) refers to the energy of the renormalised quantities, and measures the distance from the reference metric $\widetilde{{\bm{g}}}$.
Moreover ×$\mathbb{R}\times\mathbb{S}^{2}$ is endowed with the standard metric on the cylinder,
$\mathring{g}=\mathrm{d}t^{2}+d\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\,,$
and corresponding (unweighted) norms $W^{M,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ are defined in Section 4.1, to measure the size of the asymptotic geometric quantities at infinity.
Theorem 1.
Let $(\mathcal{R},{\bm{g}})$ be a solution to (1.1) in parabolic gauge, namely expressed in coordinates (1.9) relative to a time function $s:\mathcal{R}\to(0,\infty)$ whose level sets have topology ×$\mathbb{R}\times\mathbb{S}^{2}$ and satisfy the geometric gauge condition (1.12).
Suppose for some $\mathring{\varepsilon}>0$ and $s_{0}>0$, the initial data $(g_{0},k_{0})$ on $\Sigma_{s_{0}}$ is sufficiently close to the data induced by a Kerr de Sitter metric ${\bm{g}}_{\mathcal{K}_{a,m}}$ expressed in this gauge, with parameters $(a_{1},m_{1})$ at one end, and $(a_{2},m_{2})$ at the other end, in the sense that for some $N\geq 4$, with the energy defined in (1.13):
$\mathcal{E}_{N}(s_{0})=\mathring{\varepsilon}^{2}\,.$ (1.15)
  1. (I)
    Then, for $\mathring{\varepsilon}>0$ sufficiently small, the solution is global,
    $\mathcal{R}=\bigcup_{s=s_{0}}^{\infty}\Sigma_{s}\,,\qquad\text{and}\quad\mathcal{E}_{N}(s)\leq C\>\mathcal{E}_{N}(s_{0})\qquad(s\geq s_{0})\,.$ (1.16)
  2. (II)
    Furthermore, we have the following asymptotics for the spatial part of the metric (1.9):
    $g_{ij}(s,x)=g^{\infty}_{ij}(x)e^{2Hs}+h_{ij}(s,x)\,,\qquad g^{\infty}_{ij}=\widetilde{g}^{\infty}_{ij}+\widehat{g}^{\infty}_{ij}\,,\qquad h_{ij}=\widetilde{h}_{ij}+\widehat{h}_{ij}\,,$ (1.17)
    where $\widetilde{g}^{\infty}$, $g^{\infty},h,\widetilde{h}$ are metrics on ×$\mathbb{R}\times\mathbb{S}^{2}$, with $\widetilde{g}^{\infty}_{ij},\widetilde{h}_{ij}$ induced by the reference metric, satisfying
    $\|(e^{\alpha_{1}t}+e^{-\alpha_{2}t})\widehat{g}^{\infty}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}\qquad\|(e^{\alpha_{1}t}+e^{-\alpha_{2}t})\widehat{h}\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}\,.$ (1.18)
  3. (III)
    Finally, for $N\geq 6$, the lapse function admits the asymptotic expansion
    $\Phi(s,x)=1+\Phi^{\infty}(x)e^{-2Hs}+\Psi(s,x)\,,\qquad\Phi^{\infty}=\widehat{\Phi}^{\infty}+\widetilde{\Phi}^{\infty}\,,\qquad\Psi=\widehat{\Psi}+\widetilde{\Psi}\,,$ (1.19)
    where $\widetilde{\Phi}^{\infty},\widetilde{\Psi}$ are functions induced by the reference metric, satisfying
    $\begin{split}\|(e^{\alpha_{1}t}+e^{-\alpha_{2}t})\widehat{\Phi}^{\infty}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq&\,C\mathring{\varepsilon}\,,\\ \|(e^{\alpha_{1}t}+e^{-\alpha_{2}t})\widehat{\Psi}\|_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq&\,C\mathring{\varepsilon}e^{-4Hs}\,.\end{split}$ (1.20)
Proof.
The global stability statement (I) is proven in Corollary 4.11. The precise asymptotic behavior statements (II-III) are the subject of Proposition 6.1. ∎
Figure 3: Topology of the the level sets $\Sigma_{s}$ diffeomorphic to ×$\mathbb{R}\times\mathbb{S}^{2}$.
Remark 1.1 (Exponential decay).
The exponential decay assumption is not necessary: The theorem holds with $\alpha_{1}=\alpha_{2}=0$. We have included the exponential decay assumption to show that if the perturbation decays at the level of the initial data, then exponential decay is inherited along every $\Sigma_{s}$: This is true in particular for the asymptotics (1.18), (1.20). Exponential decay in Kerr de Sitter along the cosmological horizon is well-established for small angular momentum [^10] We remark that our theorem does not rely on a smallness assumption of $a_{i}$, $i=1,2$. If the black hole exteriors $\mathcal{S}_{i}$ were proven to be stable in the whole subextremal range, then our theorem provides an immediate extension to the cosmological region. in various settings: For the linear wave equation [BH08, Dya11b, Dya11a, Mav23], quasi-linear equations [Hin16, HV16, Hin17, Mav24], and – justifying our assumption here – for the Einstein equations [HV18, Fan22a, Fan22b].
Remark 1.2 (Reference metric).
The fact that it is possible to prove global existence with a fixed reference metric $\widetilde{{\bm{g}}}$ means that on every time slice $\Sigma_{s}$, $s\geq s_{0}$, the solution tends to ${\bm{g}}_{\mathcal{K}_{a,m}}$ with the same parameters – $(a_{1},m_{1})$ towards $\iota^{+}_{1}$, and $(a_{2},m_{2})$ towards $\iota^{+}_{2}$, see Fig 2. This is markedly different from the proofs of the major black hole stability theorems [HV18, KS20, DHRT21, GKS22, KS23]: In the $\Lambda>0$ case, an iterative scheme for the linearised equations is implemented in [HV18] to make successive gauge corrections and to find the parameters $(a,m)$ of the final state; in the case $\Lambda=0$ the modulation techniques [KS22a, KS22b, She23] are used in [KS23] both to anchor the gauge, and to determine the parameters $(a,m)$ of the final state. Heuristically, the reason that the parameters of the Kerr de Sitter metric remain unchanged in the cosmological region is that they are only relevant for convergence along spacelike hypersurfaces — which at their endpoints are unaffected by the perturbation.
Remark 1.3 (Functional degrees of freedom).
The main asymptotic degrees of freedom are functional in nature: [^11] This is already the case for the wave equation [Vas10, Sch15], and gives rise to a scattering problem [Ber24]. the leading orders of the solution ${\bm{g}}$ in $s$ are not captured by a member ${\bm{g}}_{\mathcal{K}_{a,m}}$ of the Kerr de Sitter family, but given by a free function, as in (1.17).  [^12] Yet locally in the past of a given point on the conformal boundary the solution is asymptotically de Sitter. Cf. discussions of the cosmic no hair conjecture in [AR16, Sch22]; for a proof in spherical symmetry see [CNO19]. The significance of this effect for the theory of gravitational radiation has been suggested by Ashtekar et al [ABK16]. We refer in this context also to the gluing and scattering constructions [Hin21, Hin24, GVK17].
Remark 1.4 (Topology).
A significant aspect of our theorem is that the topology of the spatial slices $\Sigma_{s}$ is ×$\mathbb{R}\times\mathbb{S}^{2}$ (and thus in the conformal diagram of Figure 2 the hypersurfaces $\Sigma_{s}$ extend to the end points $\iota_{1}^{+}$ and $\iota_{2}^{+}$, see Figure 3). Indeed for compact subsets $K\subset\Sigma_{s}$, with $s\geq s_{0}$ taken sufficiently large, a range of results in the literature imply that the domain of dependence $\mathcal{D}^{+}(K)$ is contained in a perturbation of de Sitter space: [^13] For a longer discussion of the spatially compact setting and its relation to earlier work on de Sitter see [Sch22, Section 1.6]. The de Sitter solution $(\mathbb{H},h)$ can be realized as a hyperboloid $\mathbb{H}$ in $\mathbb{R}^{4+1}$, with metric $h=m\rvert_{\mathbb{H}}$ induced by the ambient Minkowski metric $m$. See also [Vas10] or [Sch21] where the embedding $\mathbb{H}\subset\mathbb{R}^{4+1}$ is used, and its geometric properties are further discussed. In particular the conformal method applies [Fri86, MVK23], and more generally, Ringström proved geodesic completeness from spatially bounded data irrespective of the topology [Rin08].
Remark 1.5 (Parabolic gauge).
The gauge (1.12) is crucial for the global existence proof in this setting and in particular leads to a foliation that correctly identifies $\mathcal{I}^{+}$. A similar problem occurs in the study of stable spacelike singularities, where related ADM gauges have been used to synchronize the singularity [RS18, FRS23, FL23]; see also [AF20]. In the setting of expanding cosmologies, related gauges have been used in [LvEUW04, Fou22, FMO24].
Remark 1.6 (Global Penrose diagram).
Since [HV18] provides the stability of the regions $\mathcal{S}_{i}$ on a domain that extends beyond both the cosmological horizon and the event horizon, the result of Hintz-Vasy can in principle also be combined with a theorem of Dafermos-Luk on the $C^{0}$-stability of the Cauchy horizon [DL17, Section 1.6]. [^14] For results in spherical symmetry with $\Lambda>0$ in this region see [CGaNS15a, CGaNS15b, CGaNS17]. The combination of all three theorems shows in particular that the Penrose diagram in Fig. 1 is dynamically stable.
Acknowledgements.
This paper resolves a problem that was first suggested by Mihalis Dafermos during Volker’s Ph.D. thesis. Over the years we have benefited from conversations with many people in the field, and we would like to thank (in alphabetical order) Spyros Alexakis, Abhay Ashtekar, Peter Hintz, Gustav Holzegel, Jonathan Luk, Hans Ringström, Jared Speck, Jacques Smulevici, Jérémie Szeftel, and András Vasy for their time and interest in this project. We would also like to thank the IHP in Paris, MATRIX in Australia, and the Mittag-Leffler Institute in Sweden for their hospitality on several occasions. G.F. gratefully acknowledges the support of the ERC starting grant 101078061 SINGinGR, under the European Union’s Horizon Europe program for research and innovation. V.S. is grateful for the support from the Humboldt Professorship of Gustav Holzegel.
2 Covariant ADM formulation of the Einstein equations in parabolic gauge
In this section we decompose a $3+1$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$ with respect to a time-function $s$ which satisfies a parabolic gauge, see (1.12). For a given reference metric $\widetilde{{\bm{g}}}$, this gauge equates the deviation of the corresponding lapse functions to the deviation in the mean curvatures of the leaves $\Sigma_{s}$ of the foliation by level sets of $s$. The Einstein equations then become a first order system for the first and second fundamental forms of $\Sigma_{s}$, coupled to a parabolic equation for the lapse function.
2.1 Preliminaries of the ADM decomposition
Given a time-function $s$, a time-like vectorfield $T$ is defined by $T^{\mu}=-{\bm{g}}^{\mu\nu}\partial_{\nu}s$. The lapse function $\Phi$ is defined $\Phi^{-2}=-{\bm{g}}(T,T)$, and the unit normal to $\Sigma_{s}$ by $N=\Phi T$. For any choice of coordinates $(x^{1},x^{2},x^{3})$ on $\Sigma_{0}$, we can assign to any point $p\in\Sigma_{s}$ the coordinates $(s,x^{1},x^{2},x^{3})$ if $p=\psi_{s}(q)$, where $\psi_{s}$ is the $1$-parameter group of diffeomorphism generated by $T$, and $q\in\Sigma_{0}$ has coordinates $(x^{1},x^{2},x^{3})$. In these coordinates, the metric takes the form: [^15] As usual, we use the summation convention and Latin indices range in $\{1,2,3\}$, while Greek indices range from 00 to $3$.
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+{g}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}.$ (2.1)
We denote the first and second fundamental forms of $\Sigma_{s}$ by $g_{s}$, and $k_{s}$, respectively, and usually suppress the subscript. They are defined by
$(g_{s})_{p}={\bm{g}}_{p}\Bigr{\rvert}_{\mathrm{T}_{p}\Sigma_{s}}\,,\qquad(k_{s})_{p}(X,Y)={\bm{g}}_{p}(\nabla_{X}N,Y)\,,\qquad X,Y\in\mathrm{T}_{p}\Sigma_{s},\quad p\in\mathcal{M}.$ (2.2)
Here $N=\Phi T$ is the unit normal, and we have a coordinate frame $E_{i}=\frac{\partial}{\partial x^{i}}$ which is Lie transported by $T=\frac{\partial}{\partial s}$: $[T,E_{i}]=0$. In the frame $(E_{0}=N,E_{1},E_{2},E_{3})$ the metric components are
${\bm{g}}_{00}=-1\qquad{\bm{g}}_{0i}=0\qquad{\bm{g}}_{ij}=g_{ij}\,.$ (2.3)
The first variation formula is
$\frac{\partial g_{ij}}{\partial s}=2\Phi\,k_{ij}\,,$ (2.4)
and we will express similarly the second variation equation for $\partial_{s}k_{ij}$ as well as the Gauss-Codazzi equations of the embedding of $\Sigma_{s}$ in $\mathcal{M}$ in this frame. The derivations are well-known and can be found for instance in [Chr08].
While all spacetime quantities are set in bold, we print $\Sigma_{s}$-tangent tensors in standard font. For instance, while the Riemann curvature of $(\mathcal{M},{\bm{g}})$ is $\bm{R}_{\alpha\beta\mu\nu}$, the components of the Riemann curvature of $(\Sigma_{s},g_{s})$ are $R_{mnij}$. The Einstein vacuum equations are
${\bm{\mathrm{Ric}}}[{\bm{g}}]=\Lambda{\bm{g}}\,.$ (2.5)
While we denote the Ricci curvature of ${\bm{g}}$ by ${\bm{\mathrm{Ric}}}[{\bm{g}}]$, and the Ricci curvature of $g$ by $\operatorname{Ric}[g]$, we often denote the components of the Ricci curvature simply by:
$\operatorname{Ric}_{ij}=g^{mn}R_{minj}\,.$ (2.6)
Similarly for the Levi-Civita connections of $g,{\bm{g}}$: We denote by $\nabla$ the connection induced by ${\bm{\nabla}}$ on $\Sigma_{s}$.
The second variation formula is
$\frac{\partial k_{ij}}{\partial s}=\nabla_{i}\nabla_{j}\Phi+\Phi\Bigl{\{}-{\bm{R}}_{i0j0}+k_{i}^{m}k_{mj}\Bigr{\}}\,.$ (2.7)
The aim is to obtain a closed system of evolution equations for $g_{s}$, and $k_{s}$, and for that purpose we first eliminate the curvature component ${\bm{R}}_{0i0j}$ in (2.7) using the Einstein equations (2.5).
The Codazzi equations are:
$\nabla_{i}k_{jm}-\nabla_{j}k_{im}={\bm{R}}_{m0ij}$ (2.8)
and upon contracting and using (2.5) we obtain:
$\nabla^{i}k_{ji}-\nabla_{j}\operatorname{tr}k={\bm{\mathrm{Ric}}}_{0j}[{\bm{g}}]=0$ (2.9)
The Gauss equation reads:
$R_{minj}+k_{mn}k_{ij}-k_{mj}k_{ni}={\bm{R}}_{minj}$ (2.10)
A first contraction yields
$\operatorname{Ric}_{ij}[g]+\operatorname{tr}k\>k_{ij}-k^{n}{}_{i}\,k_{nj}={\bm{R}}_{0i0j}+{\bm{\mathrm{Ric}}}_{ij}[{\bm{g}}]={\bm{R}}_{0i0j}+\Lambda g_{ij}$ (2.11)
which gives a formula for ${\bm{R}}_{i0j0}$ which we may substitute into the second variation formula (2.7):
$\frac{\partial k_{ij}}{\partial s}=\nabla_{i}\nabla_{j}\Phi-\Phi\Bigl{\{}\operatorname{Ric}_{ij}[g]+\operatorname{tr}k\,k_{ij}-2k_{i}{}^{m}k_{mj}-{\bm{\mathrm{Ric}}}_{ij}[{\bm{g}}]\Bigr{\}}$ (2.12)
A second contraction of (2.11) gives:
$\operatorname{tr}_{g}\operatorname{Ric}+(\operatorname{tr}k)^{2}-\lvert k\rvert^{2}=2{\bm{\mathrm{Ric}}}_{00}[{\bm{g}}]+{\bm{\mathrm{tr}}}_{{\bm{g}}}{\bm{\mathrm{Ric}}}=2\Lambda$ (2.13)
where $R=\operatorname{tr}_{g}\operatorname{Ric}$ is the scalar curvature of $g$. This is the Hamiltonian constraint:
[…]
Together with (2.9), which is also referred to as the momentum constraint
$\mbox{div}_{g}k-\mathrm{d}\operatorname{tr}_{g}k=\,0\,,$ (2.15)
these are the constraint equations for the first and second fundamental form, complementing the evolution equations (2.4) for $g$ and (2.12) for $k$.
For future reference we also record the formula for the Riemann curvature of $g$ in local coordinates:
$R^{a}{}_{bmn}=\partial_{m}\Gamma^{a}_{nb}-\partial_{n}\Gamma^{a}_{mb}+\Gamma^{a}_{mc}\Gamma^{c}_{nb}-\Gamma^{a}_{nc}\Gamma^{c}_{mb}$ (2.16)
where
$\Gamma_{ic}^{a}=\frac{1}{2}g^{ab}(\partial_{i}g_{cb}+\partial_{c}g_{ib}-\partial_{b}g_{ic})\,.$ (2.17)
The curvature satisfies the cyclic identity:
$R^{a}{}_{bmn}+R^{a}{}_{mnb}+R^{a}{}_{nbm}=0\,,\quad\text{where }R^{a}{}_{bmn}=g^{ac}R_{cnbm}$ (2.18)
together with the symmetries $R^{a}_{\phantom{a}bnm}=-R^{a}_{\phantom{a}bmn}\,,\ R_{bamn}=-R_{abmn}\,,$ this implies the pair symmetry:
$R_{mnab}=R_{abmn}\,.$ (2.19)
Finally, we have in local coordinates that
$\operatorname{Ric}_{mn}=R^{a}_{\phantom{a}man}=\partial_{a}\Gamma^{a}_{nm}-\partial_{n}\Gamma^{a}_{am}+\Gamma_{ac}^{a}\Gamma_{nm}^{c}-\Gamma_{nc}^{a}\Gamma_{am}^{c}\,.$ (2.20)
The analogous formulas are valid for the Riemann curvature of ${\bm{g}}$.
2.2 System of evolution equations in parabolic gauge
We have already encountered the first variation equation (2.4) for $g_{s}$, which can also be expressed as an equation for the components of $g_{s}^{-1}$:
$\partial_{s}{g}_{ij}=$  $\,2\Phi k_{ij},$ (2.21)
$\partial_{s}g^{ij}=$ $-2\Phi k^{ij},$ (2.22)
where $k^{ij}=g^{im}g^{jn}k_{mn}$. Moreover we write the second variation equation (2.7) as
$\partial_{s}k_{ij}=\,{\nabla}_{i}{\nabla}_{j}\Phi-\Phi({\operatorname{Ric}}_{ij}+k_{ij}k_{l}{}^{l}-2{k_{i}}^{l}k_{jl})+\Phi\Lambda g_{ij}\,.$ (2.23)
In fact, it is convenient to view the second fundamental form as a $(1,1)$ tensor and work with $k_{i}{}^{j}=g^{cj}k_{ic}$ instead of $k_{ij}$. Then the equations (2.21), (2.23) become:
$\partial_{s}{g}_{ij}=$  $\,2\Phi g_{ja}k_{i}{}^{a}\,,$ (2.24)
$\partial_{s}g^{ij}=$ $-2\Phi g^{ia}k_{a}{}^{j}\,,$ (2.25)
$\partial_{s}k_{i}{}^{j}+\Phi k_{l}{}^{l}k_{i}{}^{j}=$ $\,\nabla_{i}\nabla^{j}\Phi-\Phi\operatorname{Ric}_{i}{}^{j}+\Phi\Lambda\delta_{i}{}^{j}\,.$ (2.26)
In Section 2.5 below, the Ricci curvature $\operatorname{Ric}_{i}{}^{j}$ is suitably expressed in terms of the Christoffel symbols $\Gamma_{ic}^{a}$. We are led to consider, in addition to the (2.24) and (2.26), the following evolution equation for $\Gamma_{s}$:
$\partial_{s}\Gamma_{ic}^{a}=\nabla_{i}(\Phi k_{c}{}^{a})+\nabla_{c}(\Phi k_{i}{}^{a})-g^{ab}g_{cj}\nabla_{b}(\Phi k_{i}{}^{j})$ (2.27)
This is an immediate consequence of (2.24), (2.25) and the formula (2.17):
$\begin{split}\partial_{s}\Gamma^{a}_{ic}=&-2\Phi k_{d}{}^{a}\Gamma^{d}_{ic}+g^{ab}\bigl{(}\partial_{i}(\Phi k_{cb})+\partial_{c}(\Phi k_{ib})-\partial_{b}(\Phi k_{ic})\bigr{)}\\ =&\,g^{ab}\bigl{(}\nabla_{i}(\Phi k_{cb})+\nabla_{c}(\Phi k_{ib})-\nabla_{b}(\Phi k_{ic})\bigr{)}\,.\end{split}$ (2.28)
2.2.1 Reference metric and gauge
To set up the stability problem, we consider a reference metric:
$\widetilde{{\bm{g}}}=-\widetilde{\Phi}^{2}\mathrm{d}s^{2}+\widetilde{g}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j},$ (2.29)
defined on the same differentiable manifold $\mathcal{M}$, and denote by $\widetilde{{\bm{\mathrm{Ric}}}}_{\mu\nu}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{\mu\nu}$ the components of the Ricci curvature of $\widetilde{{\bm{g}}}$. Also, we denote by ~$\widetilde{\nabla},\widetilde{\Gamma}_{ij}^{a},\widetilde{\operatorname{Ric}}_{ij}=\operatorname{Ric}[\widetilde{g}]_{ij}$ the Levi-Civita connection, Christoffel symbols, and Ricci curvature associated to $\widetilde{g}$.
Define
$\begin{split}\widehat{\Phi}=\Phi-\widetilde{\Phi},\qquad\widehat{g}_{ij}=g_{ij}-\widetilde{g}_{ij},\qquad\widehat{g}^{ij}=g^{ij}-\widetilde{g}^{ij},\\ \widehat{\nabla}=\nabla-\widetilde{\nabla},\qquad\widehat{\Gamma}_{ij}^{a}=\Gamma_{ij}^{a}-\widetilde{\Gamma}_{ij}^{a},\qquad\widehat{k}_{i}{}^{j}=k_{i}{}^{j}-\widetilde{k}_{i}{}^{j},\end{split}$ (2.30)
where
$\widetilde{k}_{i}{}^{j}=\widetilde{g}^{ja}\widetilde{k}_{ia}=\frac{1}{2}\widetilde{\Phi}^{-1}\widetilde{g}^{ja}\partial_{s}\widetilde{g}_{ia}\,.$ (2.31)
Remark 2.1.
The hats in (2.30) do not commute with the metric. For example, $g_{aj}\widehat{k}_{i}{}^{j}=\widehat{k}_{ij}\neq k_{ij}-\widetilde{k}_{ij}$, since the raising/lowering of indices of the tilde variables is performed with respect to $\widetilde{g}$. To avoid confusion, we will not change the type of the tensors with hats, that is to say, we will always treat $\widehat{k}$ as a $(1,1)$ tensor, $\widehat{g}$ as a $(0,2)$ etc.
The main remaining gauge freedom is the choice of the time-function $s$. This is the choice of a lapse function, and in this work we set
$\Phi-\widetilde{\Phi}=k_{l}{}^{l}-\widetilde{k}_{l}{}^{l}$ (2.32)
or equivalently
$\widehat{\Phi}=\widehat{k}_{l}{}^{l}\,.$ (2.33)
Remark 2.2.
A maximal gauge, where each level set of the time function has zero mean curvature,
$\operatorname{tr}_{g}k=0$ (2.34)
leads to an elliptic equation for the lapse function. Here, the choice (2.33) leads to a parabolic equation for the lapse, which is well-posed in the future direction; see (2.45) below.
Remark 2.3.
Recall that the difference of Christoffel symbols is a $(1,2)$ tensor:
$\widehat{\Gamma}_{ij}^{a}=\frac{1}{2}\widetilde{g}^{ab}(\nabla_{i}\widehat{g}_{jb}+\nabla_{j}\widehat{g}_{ib}-\nabla_{b}\widehat{g}_{ij})$ (2.35)
Also, note that
$\widetilde{g}_{ac}\widehat{\Gamma}_{ij}^{c}+\widetilde{g}_{jc}\widehat{\Gamma}_{ia}^{c}=\nabla_{i}\widehat{g}_{ja}\,.$ (2.36)
2.2.2 First and second variation equations for differences
Given a reference metric, we first derive the first variation equations for the differences $\widehat{g}_{ij}$.
Lemma 2.4 (First variation equations).
The variables $\widehat{g}_{ij},\widehat{g}^{ij}$ satisfy the evolution equations:
$\partial_{s}\widehat{g}_{ij}-2H\widehat{g}_{ij}=$  $\,2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}$ (2.37)
$\partial_{s}\widehat{g}^{ij}+2H\widehat{g}^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}-2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\,,$ (2.38)
where $H=\sqrt{\frac{\Lambda}{3}}$. Moreover
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\widehat{k}_{i}{}^{j}+\mathfrak{G}_{ic}^{a}$ (2.39)
$\mathfrak{G}_{ic}^{a}=$ $\,\Phi\widehat{\nabla}_{i}\widetilde{k}_{c}{}^{a}+\Phi\widehat{\nabla}_{c}\widetilde{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\widehat{\nabla}_{b}\widetilde{k}_{i}{}^{j}$ (2.40)
$+\widehat{\Phi}\widetilde{\nabla}_{i}\widetilde{k}_{c}{}^{a}+\widehat{\Phi}\widetilde{\nabla}_{c}\widetilde{k}_{i}{}^{a}-(\widehat{g}^{ab}g_{cj}\Phi+\widetilde{g}^{ab}\widehat{g}_{cj}\Phi+\widetilde{g}^{ab}\widetilde{g}_{cj}\widehat{\Phi})\widetilde{\nabla}_{b}\widetilde{k}_{i}{}^{j}$
$+k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi}$
$+\widehat{k}_{c}{}^{a}\widetilde{\nabla}_{i}\widetilde{\Phi}+\widehat{k}_{i}{}^{a}\widetilde{\nabla}_{c}\widetilde{\Phi}-(\widehat{g}^{ab}g_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widehat{g}_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widetilde{g}_{cj}\widehat{k}_{i}{}^{j})\widetilde{\nabla}_{b}\widetilde{\Phi}$
Proof.
To derive the equations (2.37), (2.38), (2.39), we use the fact that the corresponding variables of the reference metric satisfy the equations (2.24), (2.25), (2.27), and substract them from the equations satisfied by $g_{ij},g^{ij},\Gamma_{ic}^{a}$. The computations are straightforward. ∎
The following derivation gives the second variation equation for $\widehat{k}_{i}{}^{j}$ and shows that (2.33) is a parabolic gauge. The derivation uses a specific expression for the Ricci curvature, which we present first.
Lemma 2.5 (Ricci curvature).
The Ricci curvature of $g$ can be expressed in the form
$\begin{split}\operatorname{Ric}_{i}{}^{j}=&\frac{1}{3}g^{cj}(\nabla_{a}\Gamma_{ci}^{a}-\nabla_{c}\Gamma_{ia}^{a})+\frac{2}{3}g^{ab}(\nabla_{i}\Gamma^{j}_{ab}-\nabla_{a}\Gamma^{j}_{bi})\\ &+\frac{1}{3}g^{cj}(\Gamma^{a}_{cb}\Gamma^{b}_{ai}-\Gamma^{a}_{ab}\Gamma_{ci}^{b})+\frac{2}{3}g^{ab}(\Gamma^{c}_{ib}\Gamma_{ac}^{j}-\Gamma^{c}_{ab}\Gamma^{j}_{ci})\,.\end{split}$ (2.41)
Proof.
Starting from the expression (2.20) we can expand the expression for the Ricci curvature:
$\operatorname{Ric}_{i}{}^{j}=$ $\,g^{cj}(\partial_{a}\Gamma^{a}_{ci}-\partial_{c}\Gamma^{a}_{ia}+\Gamma^{a}_{ab}\Gamma_{ci}^{b}-\Gamma^{a}_{cb}\Gamma^{b}_{ai})$ (2.42)
$=$ $\,g^{cj}(\nabla_{a}\Gamma_{ci}^{a}-\nabla_{c}\Gamma_{ia}^{a}+\Gamma^{a}_{cb}\Gamma^{b}_{ai}-\Gamma^{a}_{ab}\Gamma_{ci}^{b}),$
where $\nabla\Gamma$ is interpreted tensorially, e.g.,
$\nabla_{a}\Gamma_{ji}^{a}:=\partial_{a}\Gamma_{ji}^{a}+\Gamma^{a}_{ab}\Gamma^{b}_{ij}-\Gamma_{aj}^{b}\Gamma_{bi}^{a}-\Gamma_{ai}^{b}\Gamma_{jb}^{a}.$
Alternatively, we write using the pair symmetry of the curvature tensor
$\operatorname{Ric}_{i}{}^{j}=$ $R^{b\phantom{ib}j}_{\phantom{b}ib}=g^{ba}R_{aib}^{\phantom{aib}j}=g^{ab}R_{b\phantom{j}ai}^{\phantom{b}j}$
$=$ $g^{ab}\bigl{(}\partial_{i}\Gamma_{ab}^{j}-\partial_{a}\Gamma_{ib}^{j}+\Gamma_{ic}^{j}\Gamma^{c}_{ab}-\Gamma_{ac}^{j}\Gamma^{c}_{ib}\bigr{)}$ (2.43)
$=$ $\,g^{ab}(\nabla_{i}\Gamma^{j}_{ab}-\nabla_{a}\Gamma^{j}_{bi}+\Gamma^{c}_{ib}\Gamma_{ac}^{j}-\Gamma^{c}_{ab}\Gamma^{j}_{ci})$
Combining (2.42) and (2.2.2) gives (2.41). ∎
The motivation for these manipulations will be discussed in Remark 2.8. We now return to (2.26).
Lemma 2.6 (Second variation equations).
The variables $\widehat{\Phi},\widehat{k}_{i}{}^{j}$ satisfy the evolution equations:
$\partial_{s}\widehat{k}_{i}{}^{j}+3H\widehat{k}_{i}{}^{j}=$ $\,g^{cj}\nabla_{i}\nabla_{c}\widehat{\Phi}-\widetilde{\Phi}(\widetilde{k}_{l}{}^{l}-3H\widetilde{\Phi}^{-1})\widehat{k}_{i}{}^{j}+\mathfrak{K}_{i}{}^{j}+(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}$ (2.44)
$+\frac{1}{3}\Phi g^{cj}(\nabla_{c}\widehat{\Gamma}_{ia}^{a}-\nabla_{a}\widehat{\Gamma}_{ci}^{a})+\frac{2}{3}\Phi g^{ab}(\nabla_{a}\widehat{\Gamma}^{j}_{bi}-\nabla_{i}\widehat{\Gamma}^{j}_{ab})$
and
$\partial_{s}\widehat{\Phi}-\Delta_{g}\widehat{\Phi}+2H\widehat{\Phi}=\mathfrak{F}+\widetilde{\mathfrak{I}}_{\Phi}$ (2.45)
where $H=\sqrt{\frac{\Lambda}{3}}$, and
$\mathfrak{K}_{i}{}^{j}=$ $-\widehat{\Phi}k_{l}{}^{l}k_{i}{}^{j}-\widetilde{\Phi}\widehat{\Phi}k_{i}{}^{j}+\Lambda\delta_{i}{}^{j}\widehat{\Phi}+\widehat{g}^{cj}\partial_{i}\partial_{c}\widetilde{\Phi}-\widehat{g}^{cj}\widetilde{\Gamma}_{ic}^{a}\partial_{a}\widetilde{\Phi}-g^{cj}\widehat{\Gamma}_{ic}^{a}\partial_{a}\widetilde{\Phi}$ (2.46)
$+\frac{1}{3}\Phi g^{cj}(\widehat{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a}-\widehat{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a})+\frac{2}{3}\Phi g^{ab}(\widehat{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi}-\widehat{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab})$
$+\frac{1}{3}(\widehat{\Phi}g^{cj}+\widetilde{\Phi}\widehat{g}^{cj})(\widetilde{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a}-\widetilde{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a})+\frac{2}{3}(\widehat{\Phi}g^{ab}+\widetilde{\Phi}\widehat{g}^{ab})(\widetilde{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi}-\widetilde{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab})$
$+\frac{1}{3}\Phi g^{cj}(\widehat{\Gamma}^{a}_{ab}\Gamma_{ci}^{b}-\widehat{\Gamma}^{a}_{cb}\Gamma^{b}_{ai})+\frac{1}{3}\Phi g^{cj}(\widetilde{\Gamma}^{a}_{ab}\widehat{\Gamma}_{ci}^{b}-\widetilde{\Gamma}^{a}_{cb}\widehat{\Gamma}^{b}_{ai})$
$+\frac{1}{3}(\widehat{\Phi}g^{cj}+\widetilde{\Phi}\widehat{g}^{cj})(\widetilde{\Gamma}^{a}_{ab}\widetilde{\Gamma}_{ci}^{b}-\widetilde{\Gamma}^{a}_{cb}\widetilde{\Gamma}^{b}_{ai})+\frac{2}{3}\Phi g^{ab}(\widehat{\Gamma}^{c}_{ab}\Gamma^{j}_{ci}-\widehat{\Gamma}^{c}_{ib}\Gamma_{ac}^{j})$
$+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j})+\frac{2}{3}(\widehat{\Phi}g^{ab}+\widetilde{\Phi}\widehat{g}^{ab})(\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j})$
$\mathfrak{F}=$ $-2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Phi\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}-\widehat{\Phi}(\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Lambda)-\widetilde{\Phi}\widehat{k}_{i}{}^{j}(\widetilde{k}_{j}{}^{i}-H\widetilde{\Phi}^{-1}\delta_{j}{}^{i})$ (2.47)
$+\widehat{g}^{ab}\partial_{a}\partial_{b}\widetilde{\Phi}-g^{ab}\widehat{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}-\widehat{g}^{ab}\widetilde{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}$
The terms (~$(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j},\widetilde{\mathfrak{I}}_{\Phi}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
(~$(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}=-\widetilde{\Phi}\Bigl{(}\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}\Bigr{)},\qquad\widetilde{\mathfrak{J}}_{\Phi}=\widetilde{\Phi}\Bigl{(}\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\Lambda\Bigr{)}\,.$ (2.48)
Proof.
For (2.44), recall that $\widetilde{{\bm{g}}}$ is not an exact solution of the Einstein equations. The starting point here is (2.12), which we can write as:
$\begin{split}\partial_{s}\widetilde{k}_{i}{}^{j}+\widetilde{\Phi}\widetilde{k}_{l}{}^{l}\widetilde{k}_{i}{}^{j}=&\,\widetilde{g}^{cj}\widetilde{\nabla}_{i}\widetilde{\nabla}_{c}\widetilde{\Phi}-\widetilde{\Phi}\operatorname{Ric}[\widetilde{g}]_{i}{}^{j}+\widetilde{\Phi}{\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{i}{}^{j}\\ =&\widetilde{g}^{cj}\widetilde{\nabla}_{i}\widetilde{\nabla}_{c}\widetilde{\Phi}-\widetilde{\Phi}\widetilde{\operatorname{Ric}}_{i}{}^{j}+\widetilde{\Phi}\Lambda\delta_{i}{}^{j}-(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\end{split}$ (2.49)
Subtracting it from (2.26) and using (2.33) then results in
$\partial_{s}\widehat{k}_{i}{}^{j}+3H\widehat{k}_{i}{}^{j}+\widehat{\Phi}k_{l}{}^{l}k_{i}{}^{j}+\widetilde{\Phi}\widehat{\Phi}k_{i}{}^{j}+\widetilde{\Phi}\bigl{(}\widetilde{k}_{l}{}^{l}-3H\widetilde{\Phi}^{-1}\bigr{)}\widehat{k}_{i}{}^{j}=\\ =\nabla_{i}\nabla^{j}\widehat{\Phi}+\widehat{g}^{jc}\widetilde{\nabla}_{i}\partial_{c}\widetilde{\Phi}+g^{jc}\widehat{\Gamma}_{ic}^{b}\partial_{b}\widetilde{\Phi}-\Phi\operatorname{Ric}_{i}{}^{j}+\widetilde{\Phi}\widetilde{\operatorname{Ric}}_{i}{}^{j}+\widehat{\Phi}\Lambda\delta_{i}{}^{j}+(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\,.$ (2.50)
This already accounts for all terms in the first line of (2.44) together with the first line in (2.46). It remains to compute the difference of the Ricci curvatures. In view of (2.41) we have:
$\begin{split}\Phi\operatorname{Ric}_{i}{}^{j}-\widetilde{\Phi}\widetilde{\operatorname{Ric}}_{i}{}^{j}=&\phantom{+}\frac{1}{3}\Phi g^{cj}(\nabla_{a}\widehat{\Gamma}_{ci}^{a}-\nabla_{c}\widehat{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\nabla_{i}\widehat{\Gamma}^{j}_{ab}-\nabla_{a}\widehat{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widehat{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\Phi g^{ab}(\widehat{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widehat{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\Phi g^{cj}(\widehat{\Gamma}^{a}_{cb}\Gamma^{b}_{ai}-\widehat{\Gamma}^{a}_{ab}\Gamma_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widehat{\Gamma}^{c}_{ib}\Gamma_{ac}^{j}-\widehat{\Gamma}^{c}_{ab}\Gamma^{j}_{ci})\\ &+\frac{1}{3}\Phi g^{cj}(\widetilde{\Gamma}^{a}_{cb}\widehat{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widehat{\Gamma}_{ci}^{b})+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci})\\ &+\frac{1}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr{)}(\widetilde{\nabla}_{a}\widetilde{\Gamma}_{ci}^{a}-\widetilde{\nabla}_{c}\widetilde{\Gamma}_{ia}^{a})+\frac{2}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr{)}(\widetilde{\nabla}_{i}\widetilde{\Gamma}^{j}_{ab}-\widetilde{\nabla}_{a}\widetilde{\Gamma}^{j}_{bi})\\ &+\frac{1}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{cj}+\Phi\widehat{g}^{cj}\bigr{)}(\widetilde{\Gamma}^{a}_{cb}\widetilde{\Gamma}^{b}_{ai}-\widetilde{\Gamma}^{a}_{ab}\widetilde{\Gamma}_{ci}^{b})+\frac{2}{3}\bigl{(}\widehat{\Phi}\widetilde{g}^{ab}+\Phi\widehat{g}^{ab}\bigr{)}(\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j}-\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci})\,.\end{split}$ (2.51)
For the equation (2.45), we first consider the contracted second variation equation, obtained by contracting (2.26):
$\partial_{s}k_{l}{}^{l}+\Phi(k_{l}{}^{l})^{2}=\Delta_{g}\phi-\Phi R+3\Lambda\Phi$ (2.52)
and eliminate the scalar curvature using the Hamiltonian constraint (2.14):
$\partial_{s}k_{l}{}^{l}=\Delta_{g}\Phi+\Lambda\Phi-\Phi|k|^{2}$ (2.53)
The corresponding equation for the mean curvature of the reference metric is found by contracting (2.49):
$\partial_{s}\widetilde{k}_{l}{}^{l}+\widetilde{\Phi}\bigl{(}\widetilde{k}_{l}{}^{l}\bigr{)}^{2}=\widetilde{\Delta}\widetilde{\Phi}-\widetilde{\Phi}\widetilde{R}+3\Lambda\widetilde{\Phi}-(\widetilde{\mathfrak{J}}_{k})_{l}{}^{l}$ (2.54)
and using the twice contracted Gauss equation:
$\widetilde{R}+(k_{l}{}^{l})^{2}-\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}=2\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\widetilde{{\bm{R}}}=2\Lambda+\widetilde{\Phi}^{-1}\widetilde{\mathfrak{J}}_{\Phi}-\widetilde{\Phi}^{-1}(\widetilde{\mathfrak{J}}_{k})_{j}{}^{j}$ (2.55)
Thus we have:
$\partial_{s}\widetilde{k}_{l}{}^{l}=\widetilde{\Delta}\widetilde{\Phi}-\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\Lambda\widetilde{\Phi}-\widetilde{\mathfrak{J}}_{\Phi}$ (2.56)
Subtracting (2.56) from (2.53) gives
$\partial_{s}\widehat{k}_{l}{}^{l}=\Delta\Phi-\widetilde{\Delta}\widetilde{\Phi}-\Phi k_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\Lambda\widehat{\Phi}+\widetilde{\mathfrak{J}}_{\Phi}$ (2.57)
Since
$\begin{split}\Phi k_{i}{}^{j}k_{j}{}^{i}-\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}=&\,\widehat{\Phi}k_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widehat{k}_{i}{}^{j}k_{j}{}^{i}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}\\ =&\,\widehat{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widehat{\Phi}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}\\ =&\,\Lambda\widehat{\Phi}+2H\widehat{k}_{j}{}^{j}\\ &+\Phi\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widehat{\Phi}\Bigl{(}\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Lambda\Bigr{)}+2\widetilde{\Phi}\widehat{k}_{i}{}^{j}\Bigl{(}\widetilde{k}_{j}{}^{i}-H\widetilde{\Phi}^{-1}\delta_{j}{}^{i}\Bigr{)}\end{split}$ (2.58)
and also
$\begin{split}\Delta\Phi-\widetilde{\Delta}\widetilde{\Phi}=&g^{ij}\nabla_{i}\partial_{j}\Phi+\widehat{g}^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}-g^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}\\ =&\Delta\widehat{\Phi}-g^{ij}\widehat{\Gamma}_{ij}^{k}\partial_{k}\widetilde{\Phi}+\widehat{g}^{ij}\widetilde{\nabla}_{i}\partial_{j}\widetilde{\Phi}\end{split}$ (2.59)
we obtain that
$\partial_{s}\widehat{k}_{l}{}^{l}+2H\widehat{k}_{l}{}^{l}=\Delta\widehat{\Phi}+\mathfrak{F}+\widetilde{\mathfrak{J}}_{\Phi}$ (2.60)
which finally implies (2.45), by virtue of the gauge condition (2.33). ∎
Finally we turn to the constraint equations for differences.
Lemma 2.7 (Constraint equations).
$\nabla_{j}\widehat{k}_{i}{}^{j}=$ $\,\partial_{i}\widehat{\Phi}-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}_{i}\,,$ (2.61)
$g^{im}\nabla_{m}\widehat{k}_{i}{}^{j}=$ $\,g^{jc}\partial_{c}\widehat{\Phi}+\widehat{g}^{jc}\widetilde{\nabla}_{c}(\widetilde{k}_{l}{}^{l}-3H)-\widehat{g}^{im}\widetilde{\nabla}_{m}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})$ (2.62)
$-g^{im}\widehat{\Gamma}_{mc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+g^{im}\widehat{\Gamma}_{mi}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}^{j}$
The terms ~$\widetilde{\mathfrak{C}}_{i},\widetilde{\mathfrak{C}}^{j}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
~$\widetilde{\mathfrak{C}}_{i}=-\widetilde{{\bm{\mathrm{Ric}}}}_{0i}\,,\qquad\widetilde{\mathfrak{C}}^{j}=\widetilde{g}^{ij}\widetilde{\mathfrak{C}}_{i}\,.$ (2.63)
Proof.
From the Codazzi equations (2.9) we know that
~$\widetilde{\nabla}_{j}\widetilde{k}_{i}{}^{j}-\widetilde{\nabla}_{i}\widetilde{k}_{l}{}^{l}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{0i}$ (2.64)
which we subtract from the momentum constraint (2.61) to get
$\nabla_{j}\widehat{k}_{i}{}^{j}-\nabla_{i}\widehat{k}_{l}{}^{l}=-\widehat{\nabla}_{j}\widetilde{k}_{i}{}^{j}+\widetilde{\mathfrak{C}}_{i}=-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}_{i}\,.$ (2.65)
In view of the gauge condition (2.33) this is (2.61). Alternatively, we can also write (2.9) as
$\widetilde{g}^{im}\widetilde{\nabla}_{m}\widetilde{k}_{i}{}^{j}-\widetilde{g}^{jc}\widetilde{\nabla}_{c}\widetilde{k}_{l}{}^{l}=\widetilde{g}^{jc}\widetilde{{\bm{\mathrm{Ric}}}}_{0c}\,,$ (2.66)
to obtain after subtracting that
$g^{im}\nabla_{m}\widehat{k}_{i}{}^{j}-g^{jc}\nabla_{c}\widehat{k}_{l}{}^{l}=-g^{im}\widehat{\nabla}_{m}\widetilde{k}_{i}{}^{j}-\widehat{g}^{im}\widetilde{\nabla}_{m}\widetilde{k}_{i}{}^{j}+\widehat{g}^{jc}\widetilde{\nabla}_{c}\widetilde{k}_{l}{}^{l}+\widetilde{\mathfrak{C}}^{j}\,.$ (2.67)
In view of the gauge condition, this gives (2.62) after expanding the first term on the RHS. ∎
Remark 2.8.
The equations (2.44), (2.39) are not symmetric hyperbolic in $\widehat{k}_{i}{}^{j},\widehat{\Gamma}_{ic}^{a}$, due to the presence of the terms $\frac{1}{3}\Phi g^{cj}\nabla_{c}\widehat{\Gamma}_{ia}^{a}$, $-\frac{2}{3}\Phi g^{ab}\nabla_{i}\widehat{\Gamma}_{ab}^{j}$ in the RHS of (2.44). However, the latter terms can be treated in the energy estimates by integrating by parts and using the constraint equations (2.61), see Section 5.1.
3 The background reference metric
In Section 2.1, we have considered a general spacetime $(\mathcal{M},{\bm{g}})$ foliated by the level sets of a time function $s$. We have also introduced coordinates $(s,x)$. In these coordinates, we will now consider a class of reference metrics of the form (2.29) which are constructed from the family of Kerr de Sitter metrics.
Definition 3.1.
We write
$f=\mathcal{O}(\eta e^{mHs})$
for some $m\in\mathbb{Z}$ and $\eta>0$, if $f(s,x)$ is a smooth (analytic) function depending only on the Kerr de Sitter metrics considered, with the property that
$|\partial_{s}^{i}\partial^{\alpha}_{x}f|\leq C_{i,\alpha}\eta e^{mHs},$ (3.1)
for any $i,\alpha$ and $(s,x)\in\mathcal{M}$.
3.1 Kerr de Sitter metric
In Boyer Lindquist coordinates $(t,r,\theta,\phi)$, the Kerr de Sitter metric reads
${\bm{g}}_{\mathcal{K}_{a,m}}=\frac{\rho^{2}}{\Delta_{r}}\mathrm{d}r^{2}+\frac{\rho^{2}}{\Delta_{\theta}}\mathrm{d}\theta^{2}+\sin^{2}\theta\frac{\Delta_{\theta}}{\rho^{2}}\bigl{(}a\mathrm{d}t-\frac{r^{2}+a^{2}}{\Delta_{0}}\mathrm{d}\phi\bigr{)}^{2}-\frac{\Delta_{r}}{\rho^{2}}\bigl{(}\mathrm{d}t-\frac{a\sin^{2}\theta}{\Delta_{0}}\mathrm{d}\phi\bigr{)}^{2},$ (3.2)
where we adopt the convention:
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\qquad\Delta_{r}=(r^{2}+a^{2})(1-\frac{\Lambda}{3}r^{2})-2mr,$ (3.3a)
$\Delta_{\theta}=1+\frac{\Lambda}{3}a^{2}\cos^{2}\theta,\qquad\Delta_{0}=1+\frac{\Lambda}{3}a^{2}.$ (3.3b)
The cosmological region is the domain $\Delta_{r}<0$, where $r$ is a time function. With the following reparametrization of the time function, to
$s=H^{-1}\ln r\quad\Leftrightarrow\quad r=e^{Hs}\,,\qquad H=\sqrt{\frac{\Lambda}{3}}\,.$ (3.4)
the Kerr de Sitter metric (3.2) then takes the form
${\bm{g}}_{\mathcal{K}_{a,m}}=-(\Phi_{\mathcal{K}_{a,m}})^{2}\mathrm{d}s^{2}+(g_{\mathcal{K}_{a,m}})_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j},\qquad x^{1}=t,\,x^{2}=\theta,\,x^{3}=\phi,$ (3.5)
where
$\Phi_{\mathcal{K}_{a,m}}=$  $\,1+\mathcal{O}(e^{-2Hs}),$
$(g_{\mathcal{K}_{a,m}})_{11}=$ $\,H^{2}e^{2Hs}+\mathcal{O}(1)\,,\qquad(g_{\mathcal{K}_{a,m}})_{13}=\,-\frac{a\sin^{2}\theta}{\Delta_{0}}H^{2}e^{2Hs}+\mathcal{O}(1)\,,$ (3.6)
$(g_{\mathcal{K}_{a,m}})_{22}=$ $\,\frac{e^{2Hs}}{\Delta_{\theta}}+\mathcal{O}(1)\,,\qquad(g_{\mathcal{K}_{a,m}})_{33}=\,\frac{\sin^{2}\theta\Delta_{\theta}+H^{2}a^{2}\sin^{4}\theta}{\Delta_{0}^{2}}e^{2Hs}+\mathcal{O}(1),$
Recall here Definition 3.1 for our use of the notation $\mathcal{O}(e^{mHs})$, for $m\in\mathbb{Z}$.
Remark 3.2.
In polar coordinates $(x^{2}=\theta,x^{3}=\phi)$ the metric components degenerate at the poles $\theta=0,\pi$. The expressions (3.1) can then be viewed as the leading order expressions of the metric components in the chart $|\theta-\pi/2|<\pi/3$, $\phi\in(0,2\pi)$. An atlas can be constructed from several of these charts, with the metric components taking identical form in each of them.
3.2 Partition of Kerr de Sitter
Let $(a_{1},m_{1})$, $(a_{2},m_{2})$ be two (possibly different) pairs of Kerr de Sitter parameters, which are sufficiently close to each other
$|a_{1}-a_{2}|+|m_{1}-m_{2}|<\widetilde{\varepsilon}.$ (3.7)
We define a metric $\widetilde{{\bm{g}}}$ as a smooth transition from ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}$ to ${\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$:
$\begin{split}\widetilde{{\bm{g}}}=&\,(1-\chi){\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}+\chi{\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}\\ =&\,{\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}}+\chi({\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}-{\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}})\end{split}$ (3.8)
where $\chi:\mathcal{M}\to[0,1]$ is a smooth function satisfying
$\chi(s,t,\theta,\phi)=\left\{\begin{array}[]{cc}0,&t\leq-1\\ 1,&t\geq 1\end{array}\right.,\qquad|\partial^{\alpha}\chi|\leq C_{\alpha},$ (3.11)
for any coordinate derivative and multi-index $\alpha$. Relative to the coordinates used in the previous subsection (recall Remark 3.2), we have
$\widetilde{{\bm{g}}}=-\widetilde{\Phi}^{2}\mathrm{d}s^{2}+\widetilde{g}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (3.12a)
where
$\widetilde{\Phi}^{2}=$ $\,(1-\chi)\Phi_{\mathcal{K}_{a_{1},m_{1}}}^{2}+\chi\Phi_{\mathcal{K}_{a_{2},m_{2}}}^{2}$ (3.12b)
$\widetilde{g}_{ij}=$ $\,(1-\chi)(g_{\mathcal{K}_{a_{1},m_{1}}})_{ij}+\chi(g_{\mathcal{K}_{a_{2},m_{2}}})_{ij}\,.$ (3.12c)
We call the reference metric (3.8) a partition of Kerr de Sitter. The main properties are recorded in the following Lemma.
Lemma 3.3.
The components of the reference metric (3.8) satisfy
$\widetilde{\Phi}-1=$ $\mathcal{O}(e^{-2Hs}),\qquad\widetilde{g}_{ij}=\mathcal{O}(e^{2Hs}),$ (3.13)
$\widetilde{\Gamma}_{ic}^{a}=$ $\mathcal{O}(1),\qquad\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j}=\mathcal{O}(e^{-2Hs}),$ (3.14)
where the $\mathcal{O}(e^{mHs})$ terms satisfy (3.1).
Proof.
The statement for $\widetilde{\Phi}$ follows from (3.12), since $\Phi_{\mathcal{K}_{a_{1},m_{1}}},\Phi_{\mathcal{K}_{a_{2},m_{2}}}$ have the same property, see (3.1). In fact, from (3.1) and (3.12), in the region where each coordinate chart is regular (cf. Remark 3.2), we also have
$C^{-1}e^{2Hs}\leq\widetilde{g}_{ij}\leq Ce^{2Hs},\qquad|\partial^{\alpha}\widetilde{g}_{ij}|\leq C_{\alpha}e^{2Hs}$
for all $(s,x)\in\mathcal{M}$. Then the statement for $\widetilde{\Gamma}_{ic}^{a}=\frac{1}{2}\widetilde{g}^{al}(\partial_{i}\widetilde{g}_{cl}+\partial_{c}\widetilde{g}_{il}-\partial_{l}\widetilde{g}_{ic})$ becomes obvious.
For $\widetilde{k}_{i}{}^{j}$ we compute
$2\Phi_{\mathcal{K}_{a,m}}(k_{\mathcal{K}_{a,m}})_{ij}=\partial_{s}(g_{\mathcal{K}_{a,m}})_{ij}=2H(g_{\mathcal{K}_{a,m}})_{ij}+\mathcal{O}(1)$
and hence
$\begin{split}(k_{\mathcal{K}_{a,m}})_{i}{}^{j}=&(g_{\mathcal{K}_{a,m}}^{-1})^{jm}(k_{\mathcal{K}_{a,m}})_{im}=\Phi_{\mathcal{K}_{a,m}}^{-1}H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})\\ =&H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})\,.\end{split}$ (3.15)
Therefore again
$2\widetilde{\Phi}\widetilde{k}_{ij}=\partial_{s}\widetilde{g}_{ij}=2H\widetilde{g}_{ij}+\mathcal{O}(1)$
and
$\widetilde{k}_{i}{}^{j}=\widetilde{g}^{jm}\widetilde{k}_{im}=\widetilde{\Phi}^{-1}H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})=H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})\,.$
This completes the proof of the lemma. ∎
3.3 Approximate solution
A partition of Kerr de Sitter is not a solution to the Einstein equations. However, all that is needed for our purposes is that it is an approximate solution. It turns out that the following is sufficient.
Proposition 3.4.
The partition metric (3.12) satisfies:
$\left\{\begin{array}[]{ccc}\big{|}\partial_{x}^{\alpha}\big{(}\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda\big{)}\big{|}&\leq\quad C_{\alpha}e^{-2Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\\ \big{|}\partial_{x}^{\alpha}\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}\big{|}&\leq\quad C_{\alpha}e^{-4Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\\ \big{|}\partial_{x}^{\alpha}\big{(}\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}\big{)}\big{|}&\leq\quad C_{\alpha}e^{-2Hs}\quad\text{or}\quad C_{\alpha}\widetilde{\varepsilon}\end{array}\right.,\qquad\text{for $|t|<1$},$ (3.19)
and
$\widetilde{\bm{\mathrm{Ric}}}_{00}+\Lambda=\widetilde{\bm{\mathrm{Ric}}}_{0}{}^{j}=\widetilde{\bm{\mathrm{Ric}}}_{i}{}^{j}+\Lambda\delta_{i}{}^{j}=0,\qquad\text{for $|t|\geq 1$}.$ (3.20)
Proof.
(3.20) is immediate from the definition of the partition metric (3.8), since for $|t|\geq 1$, $\widetilde{{\bm{g}}}$ coincides with one of the two Kerr de Sitter metrics ${\bm{g}}_{\mathcal{K}_{a_{1},m_{1}}},{\bm{g}}_{\mathcal{K}_{a_{2},m_{2}}}$.
The bounds (3.19) are proven in two steps. For the decay statement, the specific expression of the Kerr de Sitter metric is actually not used. Instead, we show that any metric which satisfies Lemma 3.3 has this property. On the other hand, the $\widetilde{\varepsilon}$ smallness of the relevant terms follows from the assumption (3.7) and the precise formula (3.2).
Step 1. Decay. The normal vectorfield $N=\widetilde{\Phi}^{-1}\partial_{s}$ satisfies ${\bm{\nabla}}_{N}N=\widetilde{\Phi}^{-1}\nabla\widetilde{\Phi}$ hence
${\bm{\nabla}}_{\partial_{s}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}{\nabla}\widetilde{\Phi}$
$\bm{\widetilde{\Gamma}}_{ss}^{s}=\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\qquad\bm{\widetilde{\Gamma}}_{ss}^{i}=\widetilde{\Phi}\widetilde{g}^{ij}\partial_{j}\widetilde{\Phi}=\mathcal{O}(e^{-4Hs})$
Moreover, we know ${\bm{\nabla}}_{\partial_{i}}N=k_{i}{}^{j}\partial_{j}$
${\bm{\nabla}}_{\partial_{i}}\partial_{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}\partial_{s}+\widetilde{\Phi}\widetilde{k}_{i}{}^{j}\partial_{j}$
$\bm{\widetilde{\Gamma}}_{is}^{s}=\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}=\mathcal{O}(e^{-2Hs})\qquad\bm{\widetilde{\Gamma}}_{is}^{j}=\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})$
Since ${\bm{\nabla}}_{\partial_{i}}\partial_{j}=\bm{\widetilde{\Gamma}}_{ij}^{s}\partial_{s}+\widetilde{\Gamma}_{ij}^{k}\partial_{k}$ where $\widetilde{\Gamma}_{ij}^{k}$ are the connection coefficients of $\widetilde{g}$, we compute
$\bm{\widetilde{\Gamma}}_{ij}^{s}=-\widetilde{\Phi}^{-1}\widetilde{{\bm{g}}}({\bm{\nabla}}_{\partial_{i}}\partial_{j},\widetilde{\Phi}^{-1}\partial_{s})=\widetilde{\Phi}^{-1}\widetilde{k}_{ij}=HG_{ij}(x)e^{2Hs}+\mathcal{O}(1)$
and
$\widetilde{\Gamma}_{ij}^{k}=\frac{1}{2}G^{kl}\bigl{(}\partial_{i}G_{jl}+\partial_{j}G_{il}-\partial_{l}G_{ij}\bigr{)}+\mathcal{O}(e^{-2Hs})$
Let us now compute the components of the Ricci curvature. We start with $\widetilde{{\bm{\mathrm{Ric}}}}_{00}$. In view of the expression (2.20), we compute
$\partial_{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\alpha}=$ $\mathcal{O}(e^{-2Hs})\,,\qquad\partial_{s}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,,$
$\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\gamma}=$ $\bm{\widetilde{\Gamma}}_{\alpha s}^{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{s}+\bm{\widetilde{\Gamma}}_{\alpha i}^{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{i}=\mathcal{O}(e^{-2Hs})\,,$
$\bm{\widetilde{\Gamma}}_{s\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}=$ $\bm{\widetilde{\Gamma}}_{sj}^{i}\bm{\widetilde{\Gamma}}_{is}^{j}+\mathcal{O}(e^{-2Hs})=3H^{2}+\mathcal{O}(e^{-2Hs})\,,$
and therefore
$\widetilde{{\bm{R}}}_{00}=\widetilde{\Phi}^{-2}\widetilde{{\bm{R}}}_{ss}=-\Lambda+\mathcal{O}(e^{-2Hs}),$ (3.21)
since $\widetilde{\Phi}^{-2}=1+\mathcal{O}(e^{-2Hs})$.
Now compute $\widetilde{{\bm{\mathrm{Ric}}}}_{0j}=\widetilde{\Phi}^{-1}\widetilde{{\bm{\mathrm{Ric}}}}_{sj}$:
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{js}$ $=\mathcal{O}(e^{-2Hs})\,,\qquad\partial_{j}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}=\mathcal{O}(e^{-2Hs})\,,$
$\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{js}^{\gamma}$ $=\bm{\widetilde{\Gamma}}_{ik}^{i}\bm{\widetilde{\Gamma}}_{js}^{k}+\mathcal{O}(e^{-2Hs})=H\widetilde{\Gamma}_{ij}^{i}+\mathcal{O}(e^{-2Hs})\,,$
$\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}$ $=\bm{\widetilde{\Gamma}}_{ji}^{s}\bm{\widetilde{\Gamma}}^{i}_{ss}+\bm{\widetilde{\Gamma}}_{jk}^{i}\bm{\widetilde{\Gamma}}_{is}^{k}+\mathcal{O}(e^{-2Hs})=H\widetilde{\Gamma}^{i}_{ji}+\mathcal{O}(e^{-2Hs})\,,$
and so by symmetry we have a cancellation
$\widetilde{{\bm{\mathrm{Ric}}}}_{0j}=\mathcal{O}(e^{-2Hs})\quad\Rightarrow\quad\widetilde{{\bm{\mathrm{Ric}}}}_{0}{}^{j}=\widetilde{g}^{aj}\widetilde{{\bm{\mathrm{Ric}}}}_{0a}=\mathcal{O}(e^{-4Hs})\,.$ (3.22)
So it remains to compute $\widetilde{{\bm{\mathrm{Ric}}}}_{ij}$:
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{ji}$ $=2H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)\,,\qquad\partial_{j}{\bm{\widetilde{\Gamma}}}^{\alpha}_{\alpha i}=\mathcal{O}(1)\,,$
$\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{ji}^{\gamma}$ $=\bm{\widetilde{\Gamma}}_{ks}^{k}{\bm{\widetilde{\Gamma}}}_{ji}^{s}+\mathcal{O}(1)=3H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)\,,$
$\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{\alpha i}^{\gamma}$ $=\bm{\widetilde{\Gamma}}_{jk}^{s}{\bm{\widetilde{\Gamma}}}_{si}^{k}+\bm{\widetilde{\Gamma}}_{js}^{k}{\bm{\widetilde{\Gamma}}}_{ki}^{s}+\mathcal{O}(1)=2H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)\,.$
Therefore,
$\widetilde{{\bm{\mathrm{Ric}}}}_{ij}=3H^{2}G_{ij}e^{2Hs}+\mathcal{O}(1)=\Lambda\widetilde{g}_{ij}+\mathcal{O}(1)$ (3.23)
and thus in view of (3.13),
$\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}=\widetilde{g}^{jk}\Bigl{(}\widetilde{{\bm{\mathrm{Ric}}}}_{ik}-\Lambda\widetilde{g}_{ik}\Bigr{)}=\mathcal{O}(e^{-2Hs})\,.$ (3.24)
Step 2. Smallness. Denote by ${}^{\mathcal{K}}\bm{\Gamma}_{\alpha\beta}^{\gamma},{}^{\mathcal{K}}{\bm{\mathrm{Ric}}}_{\mu\nu}$ the Christoffel symbol and Ricci curvature of the Kerr de Sitter metric with either $(a_{1},m_{1})$ or $(a_{2},m_{2})$ parameters. It follows directly from (3.2) and (3.7) that
$\bm{\widetilde{\Gamma}}_{ss}^{s}=$ $\,\widetilde{\Phi}^{-1}\partial_{s}\widetilde{\Phi}={{}^{\mathcal{K}}\bm{\Gamma}}_{ss}^{s}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ss}^{i}=$ $\,\widetilde{\Phi}\widetilde{g}^{ij}\partial_{j}\widetilde{\Phi}={}^{\mathcal{K}}\bm{\Gamma}_{ss}^{i}+e^{-2Hs}\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{is}^{s}=$ $\,\widetilde{\Phi}^{-1}\partial_{i}\widetilde{\Phi}={}^{\mathcal{K}}\bm{\Gamma}_{is}^{s}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{is}^{j}=$ $\,\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=\Phi_{\mathcal{K}_{a,m}}(k_{\mathcal{K}_{a,m}})_{i}{}^{j}+\mathcal{O}(\widetilde{\varepsilon})={}^{\mathcal{K}}\bm{\Gamma}_{is}^{j}+\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ij}^{s}=$ $\,\widetilde{\Phi}^{-1}\widetilde{k}_{ij}={}^{\mathcal{K}}\bm{\Gamma}_{ij}^{s}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})$
$\bm{\widetilde{\Gamma}}_{ij}^{k}=$ $\,{}^{\mathcal{K}}\bm{\Gamma}_{ij}^{k}+\mathcal{O}(\widetilde{\varepsilon})$
Therefore,
$\widetilde{{\bm{\mathrm{Ric}}}}_{ss}=\partial_{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\alpha}-\partial_{s}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}+\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{ss}^{\gamma}-\bm{\widetilde{\Gamma}}_{s\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}\\ ={}^{\mathcal{K}}{\bm{\mathrm{Ric}}}_{ss}+\mathcal{O}(\widetilde{\varepsilon})=\Lambda({\bm{g}}_{\mathcal{K}_{a,m}})_{ss}+\mathcal{O}(\widetilde{\varepsilon})=\Lambda\widetilde{{\bm{g}}}_{ss}+\mathcal{O}(\widetilde{\varepsilon}),$
because $\widetilde{{\bm{g}}}_{ss}=-\widetilde{\Phi}^{2}=-\Phi_{\mathcal{K}_{a,m}}^{2}+\mathcal{O}(\widetilde{\varepsilon})$. Hence, $\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\Lambda=\mathcal{O}(\widetilde{\varepsilon})$.
Similarly for $\widetilde{{\bm{\mathrm{Ric}}}}_{0j}=\Phi^{-1}\widetilde{{\bm{\mathrm{Ric}}}}_{sj}$:
$\widetilde{{\bm{\mathrm{Ric}}}}_{sj}=\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{js}-\partial_{j}\bm{\widetilde{\Gamma}}^{\alpha}_{\alpha s}+\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{js}^{\gamma}-\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}\bm{\widetilde{\Gamma}}_{\alpha s}^{\gamma}={}^{\mathcal{K}}{\bm{\mathrm{Ric}}}_{sj}+\mathcal{O}(\widetilde{\varepsilon})=\mathcal{O}(\widetilde{\varepsilon})$
where we have used that $\bm{\widetilde{\Gamma}}_{ji}^{s}\bm{\widetilde{\Gamma}}_{ss}^{i}=\mathcal{O}(\widetilde{\varepsilon})$.
It remains to compute $\widetilde{{\bm{\mathrm{Ric}}}}_{ij}$:
$\partial_{\alpha}\bm{\widetilde{\Gamma}}^{\alpha}_{ji}$ $=\partial_{\alpha}{}^{\mathcal{K}}\bm{\Gamma}^{\alpha}_{ji}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,,\qquad\partial_{j}{\bm{\widetilde{\Gamma}}}^{\alpha}_{\alpha i}=\partial_{j}{}^{\mathcal{K}}\bm{\Gamma}^{\alpha}_{\alpha i}+\mathcal{O}(\widetilde{\varepsilon})\,,$
$\bm{\widetilde{\Gamma}}_{\alpha\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{ji}^{\gamma}$ $={}^{\mathcal{K}}\bm{\Gamma}_{\alpha\gamma}^{\alpha}{}^{\mathcal{K}}\bm{\Gamma}_{ji}^{\gamma}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,,\qquad\bm{\widetilde{\Gamma}}_{j\gamma}^{\alpha}{\bm{\widetilde{\Gamma}}}_{\alpha i}^{\gamma}={}^{\mathcal{K}}\bm{\Gamma}_{j\gamma}^{\alpha}{}^{\mathcal{K}}\bm{\Gamma}_{\alpha i}^{\gamma}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})\,.$
Therefore
$\widetilde{{\bm{\mathrm{Ric}}}}_{ij}={\bm{\mathrm{Ric}}}[{\bm{g}}_{\mathcal{K}_{a,m}}]+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})=\Lambda\widetilde{{\bm{g}}}_{ij}+e^{2Hs}\mathcal{O}(\widetilde{\varepsilon})$
and thus in view of (3.19),
$\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}=\mathcal{O}(\widetilde{\varepsilon})\,.$
This completes the proof of the proposition. ∎
4 The bootstrap argument
4.1 Weighted norms and energy
For the spacetimes we consider, $\Sigma_{s}$ is diffeomorphic to ×$\mathbb{R}\times\mathbb{S}^{2}$. While $g=g_{s}$ is a metric on $\Sigma_{s}$, we endow (×$(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ with the standard metric $\mathring{g}$ on the cylinder:
$\mathring{g}=\mathrm{d}t^{2}+\mathring{\gamma},\qquad\mathring{\gamma}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}.$ (4.1)
The coordinate charts covering $\Sigma_{s}$ are denoted by $(t,\theta,\phi)$; recall that by construction these are coordinates on each $\Sigma_{s}$, $s\geq s_{0}$, and can also be identified with coordinates on a chart for ×$\mathbb{R}\times\mathbb{S}^{2}$. Given a $\Sigma_{s}$-tangent $(n,m)$ tensor $\mathcal{T}$, we define
$(\nabla^{(\ell)}\mathcal{T})^{j_{1}\ldots j_{n}}_{a_{1}\ldots a_{\ell}i_{1}\ldots i_{m}}=\nabla_{a_{1}}\ldots\nabla_{a_{\ell}}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}$
and
$|\mathcal{T}|_{g}^{2}=g^{i_{1}i_{1}^{\prime}}\ldots g^{i_{m}i_{m}^{\prime}}g_{j_{1}j_{1}^{\prime}}\ldots g_{j_{n}j_{n}^{\prime}}\mathcal{T}^{j^{\prime}_{1}\ldots j_{n}^{\prime}}_{i_{1}^{\prime}\ldots i_{m}^{\prime}}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}.$ (4.2)
We define ̊$\mathring{\nabla}^{(\ell)}\mathcal{T}$ and $|\mathcal{T}|_{\mathring{g}}^{2}$ similarly, using the covariant derivative $\mathring{\nabla}$ of $\mathring{g}$, instead of $\nabla$, and contracting indices with $\mathring{g}$ instead of $g$.
Definition 4.1.
Let $W^{M,\infty}(\Sigma_{s},g)$ be the space of $\Sigma_{s}$-tangent tensors with $M$ bounded spatial derivatives with respect to
$\|\mathcal{T}\|_{W^{M,\infty}(\Sigma_{s},g)}=\sum_{\ell\leq M}\text{ess\,sup}_{p\in\Sigma_{s}}e^{\ell Hs}|\nabla^{(\ell)}\mathcal{T}|_{g}(p)\,.$ (4.3)
In particular, $L^{\infty}(\Sigma_{s},g)=W^{0,\infty}(\Sigma_{s},g)$ with norm
$\|\mathcal{T}\|_{L^{\infty}(\Sigma_{s},g)}=\text{ess\,sup}_{p\in\Sigma_{s}}|\mathcal{T}|_{g}(p)\,.$ (4.4)
Definition 4.2.
Let $H^{M}(\Sigma_{s},g)$ be the Sobolev space of $\Sigma_{s}$-tangent tensors with $M$ square integrable spatial derivatives with respect to the weighted norm
$\|\mathcal{T}\|_{H^{M}(\Sigma_{s},g)}^{2}=\sum_{\ell\leq M}\int_{\Sigma_{s}}f^{2}(t)e^{2\ell Hs}|\nabla^{(\ell)}\mathcal{T}|^{2}_{g}\,e^{-3Hs}\mathrm{vol}_{g},$ (4.5)
where $\mathrm{vol}_{g}=\sqrt{|g|}\>\mathrm{d}t\wedge\mathrm{d}\theta\wedge\mathrm{d}\phi$ is the volume form of $(\Sigma_{s},g)$ and
$f(t)=e^{\alpha_{1}t}+e^{-\alpha_{2}t}$ (4.6)
is a weight function, for some $\alpha_{1},\alpha_{2}\geq 0$. In particular,
$\|\mathcal{T}\|^{2}_{L^{2}(\Sigma_{s},g)}=\int_{\Sigma_{s}}f^{2}(t)\>|\mathcal{T}|_{g}^{2}\>e^{-3Hs}\mathrm{vol}_{g}\,.$ (4.7)
We will also sometimes use the $W^{M,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g}),H^{M}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ norms, for $\Sigma_{s}$-tangent tensors, defined as follows:
$\|\mathcal{T}\|_{W^{M,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}=\sum_{\ell\leq M}\text{ess\,sup}_{p\in\mathbb{R}\times\mathbb{S}^{2}}|\mathring{\nabla}^{(\ell)}\mathcal{T}|_{\mathring{g}}(p)\,,$ (4.8)
$\|\mathcal{T}\|_{H^{M}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}=\sum_{\ell\leq M}\int_{\mathbb{R}\times\mathbb{S}^{2}}f^{2}(t)|\mathring{\nabla}^{(\ell)}\mathcal{T}|^{2}_{\mathring{g}}\mathrm{vol}_{\mathring{g}}$ (4.9)
Remark 4.3.
Note that in (4.3) and (4.5), each extra spatial derivative $\nabla$ comes at a cost of a weight $e^{Hs}$. Moreover in (4.5) the volume form is renormalised, to the effect that in this setting,
$C^{-1}\leq e^{-3Hs}\sqrt{|g|}\leq C\,,$ (4.10)
by virtue of the bootstrap assumptions (4.12) on $g$ below.
Remark 4.4.
The exponential rates $\alpha_{1},\alpha_{2}$ are related to the exponential decay of the perturbed solution towards the endpoints $t=\pm\infty$. They are nonnegative, and may not be equal; they can also be set to 00, when no exponential decay along the cosmological horizons is imposed initially. While the norm $\|\cdot\|_{H^{M}}$ does depend on $\alpha_{1},\alpha_{2}$, we typically supress this in the notation. To simplify notation, we often drop $(\Sigma_{s},g)$ from subscript to the norms (4.3), (4.4), (4.5), and (4.7).
Next, we define the overall energy for the variables $\widehat{g}_{ij},\widehat{g}^{ij},\widehat{k}_{i}{}^{j},\widehat{\Gamma}_{ic}^{a},\widehat{\Phi}$.
Definition 4.5.
If $N\in\mathbb{N}$ denotes the total number of derivatives we are commuting the main equations with, then let
$\begin{split}\mathcal{E}_{N}(s)=&\,\|\widehat{g}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{H^{N}(\Sigma_{s},g)}\\ &+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{H^{N}(\Sigma_{s},g)}\big{)}\end{split}$ (4.11)
Remark 4.6.
In terms of the $e^{Hs}$ weights, boundedness of the energy (4.11) is optimal for $\widehat{g},\widehat{g}^{-1},\widehat{\Gamma}$, since the corresponding Kerr de Sitter variables themselves do not behave better. However, the $e^{Hs}$ weights in the norms of $\widehat{\Phi},\widehat{k}$ in (4.11) are sub-optimal relative to the expected behavior of these variables ($e^{4Hs}$ would be optimal for both instead of $e^{3Hs},e^{2Hs}$). For technical reasons (hyperbolicity, boundedness of error terms etc.), we cannot propagate optimal estimates for all variables at the same time. Nevertheless, once we have completed our bootstrap argument (see Sections 4.2, 4.4), the precise asymptotic behavior of all components of the perturbed solution can be derived (see Section 6).
4.2 Bootstrap assumptions and basic consequences
Our bootstrap assumptions are that there exists a bootstrap time $s_{b}\in(s_{0},+\infty)$ such that the following inequalities hold:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{2Hs},\qquad\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq\varepsilon e^{-2Hs},\qquad\mathcal{E}_{N}(s)\leq\varepsilon^{2},$ (4.12)
for all $s\in[s_{0},s_{b})$, and some $N\geq 4$. Notice that such a bootstrap time exists from classical Cauchy stability.
4.3 Preliminary estimates
The bootstrap assumptions (4.12) have certain basic implications, which will be useful in deriving energy estimates below. First, we compare the norm (4.2) to the components.
Lemma 4.7.
Let $\mathcal{T}$ be a $\Sigma_{s}$-tangent $(n,m)$ tensor. Then the following inequalities hold:
$C^{-1}|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|\leq e^{(m-n)Hs}|\mathcal{T}|_{g}\leq C|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|$ (4.13)
for all $i_{1},\ldots,i_{m},j_{1},\ldots,j_{n}$ and $(s,x)\in(s_{0},s_{b})\times\Sigma_{s}$.
Proof.
The first two bounds in (4.12) imply that
$C^{-1}e^{2Hs}\leq g_{ij}\leq Ce^{2Hs},\qquad C^{-1}e^{-2Hs}\leq g^{ij}\leq Ce^{-2Hs},$
in a given regular coordinate patch, from which the desired inequalities readily follow. ∎
Next, we derive the Sobolev embedding for the weighted norm (4.5).
Lemma 4.8.
Let $\mathcal{T}$ be a $\Sigma_{s}$-tangent $(n,m)$ tensor. Then the following inequality holds:
$\|f(t)\mathcal{T}\|_{W^{M,\infty}(\Sigma_{s},g)}\leq C\|\mathcal{T}\|_{H^{M+2}(\Sigma_{s},g)}$ (4.14)
for all $s\in[s_{0},s_{b})$.
Proof.
The classical Sobolev embedding in ×$\mathbb{R}\times\mathbb{S}^{2}$ implies that
$f^{2}(t)|\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}|^{2}\leq C\sum_{\ell\leq 2}\int_{\Sigma_{s}}|\mathring{\nabla}^{(\ell)}[f(t)\mathcal{T}]|^{2}_{\mathring{g}}\,\mathrm{vol}_{\mathring{g}}\leq C\|\mathcal{T}\|_{H^{2}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2},$
since |̊$|\mathring{\nabla}^{(\ell)}f(t)|\leq Cf(t)$. Recall (4.10) and use Lemma 4.7 to deduce that
$\|f(t)\mathcal{T}\|_{L^{\infty}(\Sigma_{s},g)}^{2}\leq C\int_{\Sigma_{s}}f^{2}(t)\big{[}|\mathcal{T}|^{2}_{g}+e^{2Hs}|\nabla(\mathcal{T})|^{2}_{g}+e^{4Hs}|\nabla(\nabla(\mathcal{T}))|^{2}_{g}\big{]}e^{-3Hs}\mathrm{vol}_{g},$ (4.15)
where schematically
$\nabla(\mathcal{T})=$ $\,\nabla\mathcal{T}+\Gamma\star\mathcal{T}$
$\nabla(\nabla(\mathcal{T}))=$ $\,\nabla\nabla\mathcal{T}+\Gamma\star\nabla\mathcal{T}+\Gamma\star\Gamma\star\mathcal{T}+\nabla\Gamma\star\mathcal{T}$
By the first two bounds in the bootstrap assumptions (4.12), the correction terms satisfy
$|\Gamma\star\mathcal{T}|_{g}^{2}\leq$ $\,Ce^{-2Hs}|\mathcal{T}|^{2}_{g},$
$|\Gamma\star\nabla\mathcal{T}+\Gamma\star\Gamma\star\mathcal{T}+\nabla\Gamma\star\mathcal{T}|^{2}_{g}\leq$ $\,Ce^{-2Hs}|\nabla\mathcal{T}|^{2}_{g}+Ce^{-4Hs}|\mathcal{T}|^{2}_{g}.$
Inserting the former identities into (4.15) and using the latter bounds gives (4.14) for $M=0$. The proof for $M>0$ is the same, replacing $\mathcal{T}$ by $\nabla^{(\ell)}\mathcal{T}$, for $\ell\leq M$. ∎
An immediate consequence of the bootstrap assumptions (4.12) and the previous lemma is the following.
Lemma 4.9.
The variables $\widehat{g},\widehat{g}^{-1},\widehat{\Phi},\widehat{\Gamma},\widehat{k}$ satisfy the $W^{N-2,\infty}(\Sigma_{s},g)$ bound:
$\begin{split}&\|f(t)\widehat{g}\|_{W^{N-2,\infty}(\Sigma_{s},g)}+\|f(t)\widehat{g}^{-1}\|_{W^{N-2,\infty}(\Sigma_{s},g)}+e^{\frac{3}{2}Hs}\|f(t)\widehat{\Phi}\|_{W^{N-2,\infty}(\Sigma_{s},g)}\\ &+e^{Hs}\big{(}\|f(t)\widehat{\Gamma}\|_{W^{N-2,\infty}(\Sigma_{s},g)}+\|f(t)\widehat{k}\|_{W^{N-2,\infty}(\Sigma_{s},g)}\big{)}\leq C\varepsilon,\end{split}$
for all $s\in[s_{0},s_{b})$.
Next, we compare the norms defined relative to $g$ and $\mathring{g}$.
Lemma 4.10.
Let $\mathcal{T}$ be a $\Sigma_{s}$-tangent $(n,m)$ tensor. Then for $M\leq N-1$:
$C^{-1}\|\mathcal{T}\|_{W^{M,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq e^{(m-n)Hs}\|\mathcal{T}\|_{W^{M,\infty}(\Sigma_{s},g)}\leq C\|\mathcal{T}\|_{W^{M,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}$ (4.16)
and for $M\leq N$:
$C^{-1}\|\mathcal{T}\|_{H^{M}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq e^{(m-n)Hs}\|\mathcal{T}\|_{H^{M}(\Sigma_{s},g)}\leq C\|\mathcal{T}\|_{H^{M}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}.$ (4.17)
Proof.
It follows from the bootstrap assumptions (4.12) and Lemmas 4.7, 4.9. ∎
4.4 Global stability
The main energy estimates that we derive in Section 5.4 prove the following.
Theorem 2.
Assume that the bootstrap assumptions (4.12) are valid for some $N\geq 4$. Then the perturbed solution satisfies the energy estimate:
$\mathcal{E}_{N}(s)\leq C\mathcal{E}_{N}(s_{0})+C\int_{s_{0}}^{s}\,e^{-\frac{1}{2}H\tau}\mathcal{E}_{N}(\tau)\mathrm{d}\tau\\ +C\int_{s_{0}}^{s}e^{\frac{7}{2}H\tau}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big{\}}\mathrm{d}\tau$ (4.18)
for all $s\in[s_{0},s_{b})$.
Proof.
In view of the definition of the overall energy (4.11), this inequality follows directly from the main energy estimates in differential form in Proposition 5.10. Note that by adding up the inequalities (5.57)-(5.59), after multiplying the equation (5.58) for $\widehat{\Phi}$ by a suitably large constant, the term involving $\nabla^{(N+1)}\widehat{\Phi}$ on the RHS of (5.59) is absorbed by the positive term on the LHS of (5.58). ∎
The previous theorem, combined with a standard continuation argument, implies that
Corollary 4.11.
The perturbed solution exists in all of $[s_{0},+\infty)\times\Sigma_{s}$, satisfying the global estimate:
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{2Hs},\qquad\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs},\qquad\mathcal{E}_{N}(s)\leq C\mathring{\varepsilon}^{2},$ (4.19)
for all $s\in[s_{0},+\infty)$, where $\mathring{\varepsilon}^{2}:=\mathcal{E}_{N}(s_{0})$.
Proof.
Applying Gronwall’s inequality to (4.18) gives
$\mathcal{E}_{N}(s)\leq C\bigg{[}\mathcal{E}_{N}(s_{0})+\int^{s_{b}}_{s_{0}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big{\}}\mathrm{d}s\bigg{]}$
In view identities (2.48), (2.63) and Proposition 3.4, we have
$\mathcal{E}_{N}(s)\leq$ $\,C\bigg{[}\mathring{\varepsilon}^{2}+\int^{s_{*}}_{s_{0}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}ds$
$+\int^{s_{b}}_{s_{*}}e^{\frac{7}{2}Hs}\big{\{}\|\widetilde{\mathfrak{J}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\big{\}}ds\bigg{]}$
$\leq$ $\,C\big{[}\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}\big{]}$
Moreover, by Lemmas 4.8, 4.10 we obtain
$\|\widehat{g}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,Ce^{2Hs}\|\widehat{g}\|_{W^{2,\infty}(\Sigma_{s},g)}\leq Ce^{2Hs}\|\widehat{g}\|_{H^{4}(\Sigma_{s})}$
$\leq$ $\,Ce^{2Hs}\sqrt{\mathcal{E}_{N}(s)}\leq Ce^{2Hs}\sqrt{\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}},$
and similarly
$\|\widehat{g}^{-1}\|_{W^{2,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq Ce^{-2Hs}\sqrt{\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}}.$
Now choose the initial data sufficiently close to Kerr de Sitter to begin with, such that $C\mathring{\varepsilon}^{2}<\varepsilon^{2}/3$. Also, choose $s_{*}$ such that $e^{-\frac{1}{2}Hs_{*}}<\varepsilon^{2}/3$. Lastly, assume that the two Kerr de Sitter pairs of parameters $(a_{1},m_{1}),(a_{2},m_{2})$ are sufficiently close such that $\widetilde{\varepsilon}^{2}(s_{*}-s_{0})<\varepsilon^{2}/3$, to deduce that
$C\big{[}\mathring{\varepsilon}^{2}+\widetilde{\varepsilon}^{2}(s_{*}-s_{0})+e^{-\frac{1}{2}Hs_{*}}\big{]}<\varepsilon^{2}.$
Combining the above inequalities yields an improvement of our the bootstrap assumptions (4.12). By standard continuation criteria, we infer that the bootstrap time $s_{b}=+\infty$. Thus, the energy estimate (4.18), and therefore (4.19) that we have just derived using the (4.18), hold true for all $s\in[s_{0},+\infty)$. In particular, the perturbed solution exists globally. ∎
5 Future stability estimates
Our goal in this section is to derive the energy estimates that prove Theorem 2. We derive the higher order equations in Section 5.2, and treat the error estimates in Section 5.3. The overall energy estimate is proven in Proposition 5.10 at the end of Section 5.4. We start as an introduction with the basic energy identity in this gauge.
5.1 Discussion of the energy identity
The purpose of this section is to explain that $\mathcal{E}(s):=\mathcal{E}_{0}(s)$ as defined in (4.11), at zeroth order $N=0$ for simplicity, is a suitable energy for the system of equations presented in Section 2.2.2.
We begin with the first variation equations given in Lemma 2.4. The terms on the RHS of these equations will be collected as an error:
$\partial_{s}\widehat{g}_{ij}-2H\widehat{g}_{ij}=$ $\,(\mathrm{Error}_{\widehat{g}})_{ij}$ (5.1a)
$\partial_{s}\widehat{g}^{ij}+2H\widehat{g}^{ij}=$ $\,(\mathrm{Error}_{\widehat{g}^{-1}})^{ij}$ (5.1b)
In order to derive an estimate for $\widehat{g}$ in $\mathrm{L}^{2}(\Sigma_{s})$ we first derive an equation for
$|\widehat{g}|_{g}^{2}=g^{ii^{\prime}}g^{jj^{\prime}}\widehat{g}_{ij}\widehat{g}_{i^{\prime}j^{\prime}}\,.$ (5.2)
To prevent confusion, we point out that this cannot be abbreviated to $\widehat{g}^{ij}\widehat{g}_{ij}$ because we have defined in (2.30):
$\widehat{g}^{ij}=g^{ij}-\widetilde{g}^{ij}$ (5.3)
In the currrent setting, the first variation equation (2.21) can also be written as:
$\partial_{s}g_{ij}-2Hg_{ij}=(\mathrm{Error}_{g})_{ij}$ (5.4)
Similarly for $g^{-1}$. Indeed, it follows from (2.25) that
$\partial_{s}g^{ij}+2Hg^{ij}=$ $\,(\mathrm{Error}_{g^{-1}})^{ij}$ (5.5)
$(\mathrm{Error}_{g^{-1}})^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\Phi g^{ia}\bigl{(}\widetilde{k}_{a}{}^{j}-H\delta_{a}{}^{j}\Phi^{-1}\bigr{)}$ (5.6)
Therefore,
$\partial_{s}|\widehat{g}|_{g}^{2}=\mathrm{Error}_{g^{-1}}\star\widehat{g}\star\widehat{g}+\mathrm{Error}_{\widehat{g}}\star\widehat{g}$ (5.7)
where we have introduced the schematic notation $\star$ to denote all possible contractions of indices with $g$. In Lemma 5.4, a higher order version of this identity will be derived.
Next we derive the energy identity for $\widehat{\Phi}$, which indicates in particular at which rate $\widehat{\Phi}$ decays. Treating the terms on the RHS of Lemma 2.6 as error terms, we recall from (2.45) the equation for $\widehat{\Phi}$:
$\partial_{s}\widehat{\Phi}-\Delta_{g}\widehat{\Phi}+2H\widehat{\Phi}=\mathrm{Error}_{\widehat{\Phi}}$ (5.8)
After multiplying by $e^{4Hs}\widehat{\Phi}$, differentiating by parts, and rearranging the terms we obtain
$\frac{1}{2}\partial_{s}\bigl{(}e^{4Hs}\widehat{\Phi}^{2}\bigr{)}+e^{4Hs}|\nabla\widehat{\Phi}|^{2}_{g}=e^{4Hs}\nabla^{i}\bigl{(}\widehat{\Phi}\partial_{i}\widehat{\Phi}\bigr{)}+(\mathrm{Error}_{\widehat{\Phi}})e^{4Hs}\widehat{\Phi}$ (5.9)
The higher order version of this equation is given in (5.21).
Moreover, from Lemma 2.6 we rewrite the equations for $\widehat{\Gamma}$ and $\widehat{k}$ in the form
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g_{cj}\Phi\nabla^{a}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma}})_{ic}^{a}$ (5.10a)
$\partial_{s}\widehat{k}_{i}{}^{j}+3H\widehat{k}_{i}{}^{j}=$ $\,g^{cj}\nabla_{i}\nabla_{c}\widehat{\Phi}+\frac{1}{3}\Phi g^{cj}\bigl{(}\nabla_{c}\widehat{\Gamma}_{ia}^{a}-\nabla_{a}\widehat{\Gamma}_{ci}^{a}\bigr{)}+\frac{2}{3}\Phi g^{ab}\bigl{(}\nabla_{a}\widehat{\Gamma}_{bi}^{j}-\nabla_{i}\widehat{\Gamma}_{ab}^{j}\bigr{)}$ (5.10b)
$+(\mathrm{Error}_{\widehat{k}})_{i}{}^{j}$
As opposed to the equations for $\widehat{g}$ or $\widehat{\Phi}$, which can be discussed separately, the equations (5.10) need to be considered jointly, to uncover the hyperbolic structure in this formulation. Similarly to (5.7), we begin by deriving the equation for
$|\widehat{\Gamma}|_{g}^{2}=g^{ii^{\prime}}g^{cc^{\prime}}g_{aa^{\prime}}\widehat{\Gamma}_{ic}^{a}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}},$ (5.11)
which follows from (5.10a):
$\frac{1}{2}\partial_{s}|\widehat{\Gamma}|_{g}^{2}+H|\widehat{\Gamma}|_{g}^{2}=\boxed{2\Phi g^{cc^{\prime}}g_{aa^{\prime}}\widehat{\Gamma}_{ic}^{a}\nabla^{i}\widehat{k}_{c^{\prime}}{}^{a^{\prime}}}\boxed{-\Phi g^{ii^{\prime}}\widehat{\Gamma}_{ic}^{a}\nabla_{a}\widehat{k}_{i^{\prime}}{}^{c}}+\mathrm{Error}_{g,g^{-1}}\star\widehat{\Gamma}\star\widehat{\Gamma}+\mathrm{Error}_{\widehat{\Gamma}}\star\widehat{\Gamma}$ (5.12)
Moreover it follows from (5.10b) that:
$\frac{1}{2}\partial_{s}|\widehat{k}|^{2}_{g}+3H|\widehat{k}|_{g}^{2}=\widehat{k}_{i}{}^{j}\nabla^{i}\nabla_{j}\widehat{\Phi}+\frac{1}{3}\Phi g^{ii^{\prime}}\widehat{k}_{i}{}^{j}\bigl{(}\nabla_{j}\widehat{\Gamma}_{i^{\prime}a}^{a}\boxed{-\nabla_{a}\widehat{\Gamma}_{ji^{\prime}}^{a}}\bigr{)}\\ +\frac{2}{3}\Phi g^{ab}g^{ii^{\prime}}g_{jj^{\prime}}\widehat{k}_{i}{}^{j}\bigl{(}\boxed{\nabla_{a}\widehat{\Gamma}_{bi^{\prime}}^{j^{\prime}}}-\nabla_{i^{\prime}}\widehat{\Gamma}_{ab}^{j^{\prime}}\bigr{)}+\mathrm{Error}_{g,g^{-1}}\star\widehat{k}\star\widehat{k}+\mathrm{Error}_{\widehat{k}}\star\widehat{k}$ (5.13)
The crucial observation is the following: After multiplying the first equation by a factor of $1/3$, the boxed terms add up to a divergence. In fact,
$\frac{1}{2}\partial_{s}\Bigl{(}e^{2Hs}|\widehat{k}|_{g}^{2}+\frac{1}{3}e^{2Hs}|\widehat{\Gamma}|_{g}^{2}\Bigr{)}+2H|\widehat{k}|_{g}^{2}=\frac{2}{3}\Phi g^{ii^{\prime}}g_{jj^{\prime}}\nabla^{a}\bigl{(}e^{2Hs}\widehat{\Gamma}_{ai^{\prime}}^{j^{\prime}}\widehat{k}_{i}{}^{j}\bigr{)}-\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{a}\bigl{(}e^{2Hs}\widehat{\Gamma}^{a}_{ij}\widehat{k}_{i^{\prime}}{}^{j}\bigr{)}\\ +e^{2Hs}\widehat{k}_{i}{}^{j}\nabla^{i}\nabla_{j}\widehat{\Phi}+\frac{1}{3}\Phi e^{2Hs}g^{ii^{\prime}}\widehat{k}_{i}{}^{j}\nabla_{j}\widehat{\Gamma}_{i^{\prime}a}^{a}-\frac{2}{3}\Phi e^{2Hs}g^{ab}g_{jj^{\prime}}\widehat{k}_{i}{}^{j}\nabla^{i}\widehat{\Gamma}_{ab}^{j^{\prime}}\\ +e^{2Hs}\mathrm{Error}_{g,g^{-1}}\star(\widehat{\Gamma}\star\widehat{\Gamma}+\widehat{k}\star\widehat{k})+e^{2Hs}\mathrm{Error}_{\widehat{\Gamma}}\star\widehat{\Gamma}$ (5.14)
It remains to treat the terms in the second line of the above equation. After differentiation by parts, these terms produce divergences of $\widehat{k}$, and the constraint equations of Lemma 2.7 come into play:
$\nabla_{j}\widehat{k}_{i}{}^{j}=$ $\,\partial_{i}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k}})_{i}$ (5.15)
$\nabla^{i}\widehat{k}_{i}{}^{j}=$ $g^{jc}\partial_{c}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k}})^{j}$ (5.16)
The result is an equation of the form
$\frac{1}{2}\partial_{s}\Bigl{(}e^{2Hs}|\widehat{k}|_{g}^{2}+\frac{1}{3}e^{2Hs}|\widehat{\Gamma}|_{g}^{2}\Bigr{)}+2H|\widehat{k}|_{g}^{2}=\Phi\>\mathrm{div}_{g}(e^{2Hs}\widehat{\Gamma}\star\widehat{k})+\nabla^{i}\Bigl{(}e^{2Hs}\widehat{k}_{i}{}^{j}\partial_{j}\widehat{\Phi}\Bigr{)}\\ -e^{2Hs}|\nabla\widehat{\Phi}|_{g}^{2}+\frac{1}{3}\Phi e^{2Hs}\nabla^{i}\widehat{\Phi}\>\widehat{\Gamma}_{ia}^{a}-\frac{2}{3}\Phi e^{2Hs}g^{ab}\nabla_{j}\widehat{\Phi}\>\widehat{\Gamma}_{ab}^{j}\\ +e^{2Hs}\mathrm{Error}_{\mathrm{div}\widehat{k}}\star\nabla\widehat{\Phi}+e^{2Hs}\mathrm{Error}_{\mathrm{div}\widehat{k},\widehat{\Gamma}}\star\widehat{\Gamma}+e^{2Hs}\mathrm{Error}_{g,g^{-1}}\star(\widehat{\Gamma}\star\widehat{\Gamma}+\widehat{k}\star\widehat{k})$ (5.17)
The higher order version of this equation is the content of Lemma 5.6.
It is now clear from (5.7), (5.9), and (5.17) that a suitable energy for this system is indeed
$\mathcal{E}(s)=\,\|\widehat{g}\|^{2}_{L^{2}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{L^{2}(\Sigma_{s},g)}+e^{3Hs}\|\widehat{\Phi}\|^{2}_{L^{2}(\Sigma_{s},g)}+e^{2Hs}\big{(}\|\widehat{\Gamma}\|^{2}_{L^{2}(\Sigma_{s},g)}+\|\widehat{k}\|^{2}_{L^{2}(\Sigma_{s},g)}\big{)}$ (5.18)
Note however that in comparison to (5.9), the rate for $\widehat{\Phi}$ included in the energy is not sharp. This gap is needed to close the energy estimates; see for example the error estimates of Lemma 5.8 below. The higher order version of the energy defined above is precisely (4.11). Nevertheless, sharp asymptotics for $\widehat{\Phi}$ and the rest of the variables are derived in Section 6, after we have completed the overall energy argument.
For the energy estimates, note already that $\nabla\widehat{\Phi}$ appears on the RHS of (5.17) at one order of differentiability higher than in the energy (5.18). Here the positive term on the LHS of (5.9) is used. However, to proceed we need to estimate in the first place various errors. This will be done systematically in Section 5.3.
5.2 The differentiated equations and higher order energy identities
5.2.1 Higher order equations
To derive higher order energy estimates, we first commute the first and second variation equations of Section 2.2.2 with tangential derivatives.
Lemma 5.1 (Commuted first variation equations).
$\partial_{s}\nabla^{(\ell)}\widehat{g}_{ij}-2H\nabla^{(\ell)}\widehat{g}_{ij}=$ $\,(\mathrm{Error}_{\widehat{g},\ell})_{ij},$ (5.19a)
$\partial_{s}\nabla^{(\ell)}\widehat{g}^{ij}+2H\nabla^{(\ell)}\widehat{g}^{ij}=$ $\,(\mathrm{Error}_{\widehat{g}^{-1},\ell})^{ij}\,,$ (5.19b)
where
$(\mathrm{Error}_{\widehat{g},\ell})_{ij}=$ ()$\,\nabla^{(\ell)}\big{\{}2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}\big{\}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}_{ij},$ (5.20a)
$(\mathrm{Error}_{\widehat{g}^{-1},\ell})^{ij}=$ ()$-\nabla^{(\ell)}\big{\{}2\Phi g^{ia}\widehat{k}_{a}{}^{j}+2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}+2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\big{\}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}^{ij}\,.$ (5.20b)
Moreover,
$\partial_{s}\nabla^{(\ell)}\widehat{\Phi}-\Delta_{g}\nabla^{(\ell)}\widehat{\Phi}+2H\nabla^{(\ell)}\widehat{\Phi}=\,\mathrm{Error}_{\widehat{\Phi},\ell},$ (5.21)
where
$\mathrm{Error}_{\widehat{\Phi},\ell}=\,\nabla^{(\ell)}(\mathfrak{F}+\widetilde{\mathfrak{I}}_{\Phi})+[\partial_{s},\nabla^{(\ell)}]\widehat{\Phi}-[\Delta_{g},\nabla^{(\ell)}]\widehat{\Phi}\,.$ (5.22)
Proof.
These equations result by commuting the equations (2.37), (2.38), (2.45) with $\nabla^{(\ell)}$. ∎
Lemma 5.2 (Commuted second variation equations).
$\begin{split}\partial_{s}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}+3H\nabla^{(\ell)}\widehat{k}_{i}{}^{j}=&\,\frac{1}{3}\Phi g^{cj}(\nabla_{c}\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}-\nabla_{a}\nabla^{(\ell)}\widehat{\Gamma}_{ci}^{a})+(\mathrm{Error}_{\widehat{k},\ell})_{i}{}^{j}\\ &+\frac{2}{3}\Phi g^{ab}(\nabla_{a}\nabla^{(\ell)}\widehat{\Gamma}^{j}_{bi}-\nabla_{i}\nabla^{(\ell)}\widehat{\Gamma}^{j}_{ab})+g^{cj}\nabla_{i}\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\end{split}$ (5.23)
where
$(\mathrm{Error}_{\widehat{k},\ell})_{i}{}^{j}=$ ()$\,\nabla^{(\ell)}\big{\{}\mathfrak{K}_{i}{}^{j}-\widetilde{\Phi}(\widetilde{k}_{l}{}^{l}-3H\widetilde{\Phi}^{-1})\widehat{k}_{i}{}^{j}+(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\big{\}}$ (5.24)
$+\frac{1}{3}\Phi g^{cj}\big{(}[\nabla^{(\ell)},\nabla_{c}]\widehat{\Gamma}_{ia}^{a}-[\nabla^{(\ell)},\nabla_{a}]\widehat{\Gamma}_{ci}^{a}\big{)}+g^{cj}[\nabla^{(\ell)},\nabla_{i}]\nabla_{c}\widehat{\Phi}$
$+\frac{2}{3}\Phi g^{ab}\big{(}[\nabla^{(\ell)},\nabla_{a}]\widehat{\Gamma}^{j}_{bi}-[\nabla^{(\ell)},\nabla_{i}]\widehat{\Gamma}^{j}_{ab}\big{)}+[\partial_{s},\nabla^{(\ell)}]\widehat{k}_{i}{}^{j}$
$+\sum_{\ell_{1}+\ell_{2}=\ell,\,\ell_{2}<\ell}\big{\{}\frac{1}{3}g^{cj}\nabla^{(\ell_{1})}\Phi(\nabla^{(\ell_{2})}\nabla_{c}\widehat{\Gamma}_{ia}^{a}-\nabla^{(\ell_{2})}\nabla_{a}\widehat{\Gamma}_{ci}^{a})$
$+\frac{2}{3}g^{ab}\nabla^{(\ell_{1})}\Phi(\nabla^{(\ell_{2})}\nabla_{a}\widehat{\Gamma}^{j}_{bi}-\nabla^{(\ell_{2})}\nabla_{i}\widehat{\Gamma}^{j}_{ab})\big{\}}\,.$
Moreover
$\partial_{s}\nabla^{(\ell)}\widehat{\Gamma}_{ic}^{a}=\Phi\nabla_{i}\nabla^{(\ell)}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\nabla^{(\ell)}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}+(\mathrm{Error}_{\widehat{\Gamma},\ell})_{ic}^{a}$ (5.25)
where
$(\mathrm{Error}_{\widehat{\Gamma},\ell})_{ic}^{a}=$ $\,\nabla^{(\ell)}\mathfrak{G}_{ic}^{a}+[\partial_{s},\nabla^{(\ell)}]\widehat{\Gamma}^{a}_{ic}+\Phi[\nabla^{(\ell)},\nabla_{i}]\widehat{k}_{c}{}^{a}+\Phi[\nabla^{(\ell)},\nabla_{c}]\widehat{k}_{i}{}^{a}$
$-g^{ab}g_{cj}\Phi[\nabla^{(\ell)},\nabla_{b}]\widehat{k}_{i}{}^{j}+\sum_{\ell_{1}+\ell_{2}=\ell,\,\ell_{2}<\ell}\big{\{}\nabla^{(\ell_{1})}\Phi\nabla^{(\ell_{2})}\nabla_{i}\widehat{k}_{c}{}^{a}+\nabla^{(\ell_{1})}\Phi\nabla^{(\ell_{2})}\nabla_{c}\widehat{k}_{i}{}^{a}$ (5.26)
$-g^{ab}g_{cj}\nabla^{(\ell_{1})}\Phi\nabla^{(\ell_{2})}\nabla_{b}\widehat{k}_{i}{}^{j}\big{\}}$
Proof.
These result by commuting the equations (2.44) and (2.39) with $\nabla^{(\ell)}$. ∎
Finally, we also commute the constraint equations (2.61), (2.62) with $\nabla^{(\ell)}$.
Lemma 5.3 (Commuted constraint equations).
$\nabla_{j}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}=$ $\,\nabla_{i}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})_{i},$ (5.27a)
$g^{im}\nabla_{m}\nabla^{(\ell)}\widehat{k}_{i}{}^{j}=$ $\,g^{jc}\nabla_{c}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{j},$ (5.27b)
where
$(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})_{i}=$ (){~$\,\nabla^{(\ell)}\big{\{}\widetilde{\mathfrak{C}}_{i}-\widehat{\Gamma}_{jc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})+\widehat{\Gamma}_{ji}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})\big{\}}$ (5.28a)
$+[\nabla^{(\ell)},\nabla_{i}]\widehat{\Phi}+[\nabla_{j},\nabla^{(\ell)}]\widehat{k}_{i}{}^{j},$
$(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{j}=$ ()$\,\nabla^{(\ell)}\big{\{}\widehat{g}^{jc}\widetilde{\nabla}_{c}(\widetilde{k}_{l}{}^{l}-3H)-\widehat{g}^{im}\widetilde{\nabla}_{m}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})-g^{im}\widehat{\Gamma}_{mc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})$ (5.28b)
$+g^{im}\widehat{\Gamma}_{mi}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})+\widetilde{\mathfrak{C}}^{j}\big{\}}+g^{jc}[\nabla^{(\ell)},\nabla_{c}]\widehat{\Phi}+g^{im}[\nabla_{m},\nabla^{(\ell)}]\widehat{k}_{i}{}^{j}.$
5.2.2 Higher order energy identities
Next, we write the energy identities for the above system of equations, at the level of pointwise magnitudes $|\nabla^{\ell}\widehat{g}|_{g}^{2}$, $|\nabla^{\ell}\widehat{\Phi}|_{g}^{2}$, and so forth.
Lemma 5.4 (Energy identity for $\widehat{g}$).
$\sum_{\ell\leq N}\frac{1}{2}\partial_{s}\Bigl{\{}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}+e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}^{-1}|_{g}^{2}\Bigr{\}}\\ =\sum_{\ell\leq N}e^{2\ell Hs}\Bigl{\{}(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{g}\star\nabla^{(\ell)}\widehat{g}+(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{g}^{-1}\star\nabla^{(\ell)}\widehat{g}^{-1}\\ +\mathrm{Error}_{\widehat{g},\ell}\star\nabla^{(\ell)}\widehat{g}+\mathrm{Error}_{\widehat{g}^{-1},\ell}\star\nabla^{(\ell)}\widehat{g}^{-1}\Bigr{\}}\,.$ (5.29)
Proof.
To differentiate
$e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}=e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}$ (5.30)
we use (2.25) and (5.19a) and obtain
$\begin{split}\frac{1}{2}\partial_{s}\big{\{}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\big{\}}=&\,\ell He^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\\ &-\Phi k_{a}{}^{b_{1}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}a}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ &-\ldots-\Phi k_{a}{}^{b_{\ell}^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}a}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{i^{\prime}}e^{2\ell Hs}g^{ia}g^{jj^{\prime}}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ -&\Phi k_{a}{}^{j^{\prime}}e^{2\ell Hs}g^{ii^{\prime}}g^{ja}g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ +e^{2\ell Hs}g^{ii^{\prime}}g^{jj^{\prime}}&g^{b_{1}b_{1}^{\prime}}\ldots g^{b_{\ell}b_{\ell}^{\prime}}\big{\{}2H\nabla_{b_{1}}\ldots\nabla_{b_{\ell}}\widehat{g}_{ij}+(\mathrm{Error})_{\widehat{g},\ell})_{ij}\big{\}}\nabla_{b_{1}^{\prime}}\ldots\nabla_{b_{\ell}^{\prime}}\widehat{g}_{i^{\prime}j^{\prime}}\\ \end{split}$ (5.31)
Since
$\Phi k_{a}^{b}=\Phi\widehat{k}_{a}{}^{b}+\Phi\widetilde{k}_{a}{}^{b}-H\delta_{a}{}^{b}+H\delta_{a}{}^{b}$ (5.32)
the diagonal terms cancel and we are left with
$\frac{1}{2}\partial_{s}\big{\{}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{g}|_{g}^{2}\big{\}}=\,e^{2\ell Hs}\Bigl{\{}(\Phi\widehat{k}+\Phi\widetilde{k}-H\delta)\star\nabla^{(\ell)}\widehat{g}\star\nabla^{(\ell)}\widehat{g}+\mathrm{Error}_{\widehat{g},\ell}\star\nabla^{(\ell)}\widehat{g}\Bigr{\}}\,.$ (5.33)
Similarly for $\widehat{g}^{-1}$. ∎
Lemma 5.5 (Energy identity for $\widehat{\Phi}$).
$\sum_{\ell\leq N}\bigg{[}\frac{1}{2}\partial_{s}\big{\{}e^{3Hs}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{\Phi}|^{2}_{g}\big{\}}+e^{3Hs}e^{2\ell Hs}|\nabla^{(\ell+1)}\widehat{\Phi}|^{2}_{g}+2He^{3Hs}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{\Phi}|^{2}_{g}\bigg{]}\\ =\sum_{\ell\leq N}e^{3Hs}e^{2\ell Hs}\big{\{}\nabla^{i}\big{[}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big{]}+(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{\Phi}\star\nabla^{(\ell)}\widehat{\Phi}+\mathrm{Error}_{\widehat{\Phi},\ell}\star\nabla^{(\ell)}\widehat{\Phi}\big{\}}$ (5.34)
Proof.
We multiply equation (5.21) with $e^{3Hs}e^{2\ell Hs}\nabla^{(\ell)}\widehat{\Phi}$, differentiate by parts in $\partial_{s}$, and differentiate by parts in $\nabla_{i}$ the term with $\ell+2$ spatial derivatives; contract all corresponding pairs of indices using the metric, and sum in $\ell\leq N$. The correction term with factor $\Phi k-H\delta$ arises in the integration by parts in $\partial_{s}$ as in the proof of Lemma 5.4. ∎
Lemma 5.6 (Energy identity for $\widehat{\Gamma}$ and $\widehat{k}$).
$\sum_{\ell\leq N}\bigg{[}\frac{1}{2}\partial_{s}\big{\{}\frac{1}{3}e^{2Hs}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{\Gamma}|^{2}_{g}+e^{2Hs}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{k}|^{2}_{g}\big{\}}+2He^{2Hs}e^{2\ell Hs}|\nabla^{(\ell)}\widehat{k}|^{2}_{g}\bigg{]}$
$=\sum_{\ell\leq N}\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell}+\sum_{\ell\leq N}e^{2Hs}e^{2\ell Hs}\bigg{[}\frac{2}{3}\Phi g^{ab}g_{jj^{\prime}}\big{\{}g^{j^{\prime}c}\nabla_{c}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{j^{\prime}}\big{\}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ab}^{j}$
$-\frac{1}{3}\Phi g^{ii^{\prime}}\big{\{}\nabla_{i^{\prime}}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})_{i^{\prime}}\big{\}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}-\big{\{}\nabla^{c}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{c}\big{\}}\star\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\bigg{]}$ (5.35)
$+\sum_{\ell\leq N}e^{2Hs}e^{2\ell Hs}\big{\{}(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{k}\star\nabla^{(\ell)}\widehat{k}+(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{\Gamma}\star\nabla^{(\ell)}\widehat{\Gamma}$
$+\mathrm{Error}_{\widehat{\Gamma},\ell}\star\nabla^{(\ell)}\widehat{\Gamma}+\mathrm{Error}_{\widehat{k},\ell}\star\nabla^{(\ell)}\widehat{k}\big{\}},$
where
$\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell}=e^{2Hs}e^{2\ell Hs}\bigg{[}\frac{2}{3}\Phi g^{ii^{\prime}}g^{cc^{\prime}}g_{aa^{\prime}}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}\star\nabla^{(\ell)}\widehat{k}_{c}{}^{a}\big{]}\\ -\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{b}\big{[}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}j}^{b}\star\nabla^{(\ell)}\widehat{k}_{i}{}^{j}\big{]}+\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{c}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big{]}\\ -\frac{2}{3}\Phi g^{ab}g^{ii^{\prime}}g_{jj^{\prime}}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ab}^{j}\big{]}+g^{ii^{\prime}}g_{jj^{\prime}}g^{cj}\nabla_{i}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\star\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\big{]}\bigg{]}$ (5.36)
where the $\star$ symbol in all schematic expressions above signifies that all relevant indices in these terms are contracted.
Proof.
This energy identity follows from the higher order equations of Lemma 5.2: First multiply
(5.25)×$\times\frac{1}{3}e^{2Hs}e^{2\ell Hs}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}$
and contract all corresponding pairs of indices $(i;i^{\prime})$, $(j;j^{\prime})$, $(c;c^{\prime})$ using the metric. Similarly multiply and contract
(5.23)×$\times e^{2Hs}e^{2\ell Hs}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}$
and sum up the resulting equations. After differentiating by parts in $\partial_{s}$, we obtain the principal terms on the LHS of (5.6), and the error terms involving $\Phi k-H\delta$ on the RHS.
It remains to differentiate by parts the terms in the RHS which contain $\ell+1$ spatial derivatives of $\widehat{\Gamma}$ and $\widehat{k}$. This produces on one hand the divergence terms $(\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell})$ in (5.36), and on the other hand, divergences of $\nabla^{(\ell)}k$ which we can replace using the higher order constraint equations (5.27). For instance,
$g^{ii^{\prime}}g_{jj^{\prime}}e^{2Hs}e^{2\ell Hs}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{j^{\prime}}\frac{1}{3}\Phi g^{cj}\nabla_{c}\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}=$
$=$ $\,e^{2Hs}e^{2\ell Hs}\big{\{}\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{c}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big{]}-\frac{1}{3}\Phi g^{ii^{\prime}}(\nabla_{c}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c})\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big{\}}$
$=$ $\,e^{2Hs}e^{2\ell Hs}\bigg{[}\frac{1}{3}\Phi g^{ii^{\prime}}\nabla_{c}\big{[}\nabla^{(\ell)}\widehat{k}_{i^{\prime}}{}^{c}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\big{]}-\frac{1}{3}\Phi g^{ii^{\prime}}\big{\{}\nabla_{i^{\prime}}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})_{i^{\prime}}\big{\}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ia}^{a}\bigg{]}$
The rest of the computations are similar and straightforward. ∎
5.3 Error estimates
In this section we estimate the error terms in (5.20), (5.22), (5.24) and in (5.28), assuming the bootstrap assumptions (4.12) are valid. For this purpose, we first derive commutator estimates.
Lemma 5.7.
Let $\mathcal{T}$ be $\Sigma_{s}$-tangent $(n,m)$ tensor and let $\ell\leq N$. Then it satisfies:
$e^{\ell Hs}\|[\partial_{s},\nabla^{(\ell)}]\mathcal{T}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell-1}(\Sigma_{s},g)},$ (5.37a)
$e^{\ell Hs}\|[\nabla_{i},\nabla^{(\ell)}]\mathcal{T}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell-1}(\Sigma_{s},g)},$ (5.37b)
$e^{(\ell+1)Hs}\|[\Delta_{g},\nabla^{(\ell)}]\mathcal{T}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell}(\Sigma_{s},g)},$ (5.37c)
for all $s\in[s_{0},s_{b})$.
Proof.
First, we derive formulas for the commutators. Commuting $\partial_{s}$ with $\nabla$ applied to $\mathcal{T}$ gives:
$\partial_{s}\nabla_{b}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}=\,\partial_{s}\big{\{}\partial_{b}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}+$ $\Gamma_{bc}^{j_{1}}\mathcal{T}^{c\ldots j_{n}}_{i_{1}\ldots i_{m}}+\ldots+\Gamma_{bc}^{j_{n}}\mathcal{T}^{j_{1}\ldots c}_{i_{1}\ldots i_{m}}$
$-\Gamma_{bi_{1}}^{c}\mathcal{T}^{j_{1}\ldots j_{n}}_{c\ldots i_{m}}-\ldots-\Gamma_{bi_{m}}^{c}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots c}\big{\}}$
$=\,\nabla_{b}\partial_{s}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots i_{m}}+$ $\partial_{s}\Gamma_{bc}^{j_{1}}\mathcal{T}^{c\ldots j_{n}}_{i_{1}\ldots i_{m}}+\ldots+\partial_{s}\Gamma_{bc}^{j_{n}}\mathcal{T}^{j_{1}\ldots c}_{i_{1}\ldots i_{m}}$
$-\partial_{s}\Gamma_{bi_{1}}^{c}\mathcal{T}^{j_{1}\ldots j_{n}}_{c\ldots i_{m}}-\ldots-\partial_{s}\Gamma_{bi_{m}}^{c}\mathcal{T}^{j_{1}\ldots j_{n}}_{i_{1}\ldots c}.$
Using (2.27) to replace $\partial_{s}\Gamma$, we have schematically
$[\partial_{s},\nabla]\mathcal{T}=\nabla(\Phi k)\star\mathcal{T},$
and by induction on $\ell$:
$[\partial_{s},\nabla^{(\ell)}]\mathcal{T}=\sum_{\ell_{1}+\ell_{2}=\ell,\,\ell_{2}<\ell}\nabla^{\ell_{1}}(\Phi k)\star\nabla^{\ell_{2}}\mathcal{T}$
For the estimate (5.37a), let us make a case distinction for the terms in this sum, depending on $\ell_{1}$. For $20<\ell_{1}\leq N-$ 2 we have
$e^{\ell Hs}\|\nabla^{\ell_{1}}(\Phi k)\star\nabla^{\ell_{2}}\mathcal{T}\|_{L^{2}}\leq e^{\ell_{1}Hs}\|\nabla^{(\ell_{1})}(\Phi k)\|_{L^{\infty}}\|\mathcal{T}\|_{H^{\ell_{2}}}\leq Ce^{-Hs}\|\mathcal{T}\|_{H^{\ell-1}}$ (5.38)
where we have used Lemma 4.9, Lemma 3.3, and Lemma 4.10. For $\ell_{1}>N-2$ the estimate still holds for those terms in
$\nabla^{(\ell_{1})}\bigl{(}\Phi k\bigr{)}=\nabla^{(\ell_{1})}\bigl{(}\widehat{\Phi}\widehat{k}+\widetilde{\Phi}\widehat{k}+\widehat{\Phi}\widetilde{k}+\widetilde{\Phi}\widetilde{k}\bigr{)}$ (5.39)
involving at most $N-2$ derivatives of $\widehat{\Phi}$, and $\widehat{k}$. For the remaining terms, with more that $N-2$ derivatives of $\widehat{\Phi}$ or $\widehat{k}$, we use the bootstrap assumption (4.12) on the energy, and Lemma 4.14 for $\mathcal{T}$. For example:
$e^{\ell Hs}\|(\nabla^{(\ell_{1})}\widehat{\Phi})\widetilde{k}\star\nabla^{\ell_{2}}\mathcal{T}\|_{L^{2}}\leq\|\widehat{\Phi}\|_{H^{\ell_{1}}}\|\mathcal{T}\|_{W^{\ell_{2},\infty}}\leq Ce^{-\frac{3}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}\|\mathcal{T}\|_{H^{\ell_{2}+2}}\,,$ (5.40)
where $\ell_{1}\geq N-1\geq 3$ and $\ell_{2}+2=\ell-\ell_{1}+2\leq\ell-1$.
For the commutation of spatial derivatives we write schematically:
$[\nabla_{i},\nabla^{(\ell)}]\mathcal{T}=$ $\,\sum_{\ell_{1}+\ell_{2}=\ell,\,\ell_{2}<\ell}\nabla^{(\ell_{1})}\Gamma\star\nabla^{(\ell_{2})}\mathcal{T}+\!\!\!\!\sum_{\begin{subarray}{c}\ell_{1}+\ell_{2}+\ell_{3}=\ell-1\\ \ell_{3}<\ell\end{subarray}}\!\!\!\!\nabla^{(\ell_{1})}\Gamma\star\nabla^{(\ell_{2})}\Gamma\star\nabla^{(\ell_{3})}\mathcal{T}$ (5.41)
$[\Delta_{g},\nabla^{(\ell)}]\mathcal{T}=$ $\,\nabla^{i}[\nabla_{i},\nabla^{(\ell)}]\mathcal{T}+[\nabla^{i},\nabla^{(\ell)}]\nabla_{i}\mathcal{T}$ (5.42)
The estimates (5.37b) and (5.37c) then follow by estimating the above expressions using the bootstrap assumptions (4.12) and the Lemma 4.8. ∎
Next, we estimate the error terms coming from (2.40), (2.46), (2.47).
Lemma 5.8.
The expressions $\mathfrak{G}_{ic}^{a},\mathfrak{K}_{i}{}^{j},\mathfrak{F}$ satisfy:
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}\mathfrak{G}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{-\frac{1}{2}Hs}e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}$ (5.43a)
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}\mathfrak{K}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}$ (5.43b)
$e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell)}\mathfrak{F}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}$ (5.43c)
for all $s\in[s_{0},s_{b})$ and $\ell\leq N$.
Proof.
We begin with (5.43a). Going back to (2.40), we notice the following cancellations in the first line:
$\begin{split}\Phi\widehat{\nabla}_{i}\widetilde{k}_{c}{}^{a}&+\Phi\widehat{\nabla}_{c}\widetilde{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\widehat{\nabla}_{b}\widetilde{k}_{i}{}^{j}=\\ =&\,\Phi\widehat{\Gamma}_{ib}^{a}\widetilde{k}_{c}{}^{b}-\Phi\widehat{\Gamma}_{ic}^{b}\widetilde{k}_{b}{}^{a}+\Phi\widehat{\Gamma}_{cb}^{a}\widetilde{k}_{i}{}^{b}-\Phi\widehat{\Gamma}_{ic}^{b}\widetilde{k}_{b}{}^{a}-g^{ab}g_{cj}\Phi(\widehat{\Gamma}_{bc}^{j}\widetilde{k}_{i}{}^{c}-\widehat{\Gamma}_{bi}^{c}\widetilde{k}_{c}{}^{j})\\ =&\,\Phi\widehat{\Gamma}_{ib}^{a}(\widetilde{k}_{c}{}^{b}-H\delta_{c}{}^{b})-\Phi\widehat{\Gamma}_{ic}^{b}(\widetilde{k}_{b}{}^{a}-H\delta_{b}{}^{a})+\Phi\widehat{\Gamma}_{cb}^{a}(\widetilde{k}_{i}{}^{b}-H\delta_{i}{}^{b})-\Phi\widehat{\Gamma}_{ic}^{b}(\widetilde{k}_{b}{}^{a}-H\delta_{b}{}^{a})\\ &-g^{ab}g_{cj}\Phi\big{[}\widehat{\Gamma}_{bc}^{j}(\widetilde{k}_{i}{}^{c}-H\delta_{i}{}^{c})-\widehat{\Gamma}_{bi}^{c}(\widetilde{k}_{c}{}^{j}-H\delta_{c}{}^{j})\big{]}\end{split}$ (5.44)
Using the bootstrap assumptions (4.12) and Lemma 4.9, as well as the properties of the reference metric in Lemma 3.3, we deduce that
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}(\Phi\widehat{\nabla}_{i}\widetilde{k}_{c}{}^{a}+\Phi\widehat{\nabla}_{c}\widetilde{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\widehat{\nabla}_{b}\widetilde{k}_{i}{}^{j})\|_{L^{2}(\Sigma_{s},g)}\leq Ce^{-2Hs}\sqrt{\mathcal{E}_{N}(s)},$
which of course is much better than the asserted bound for $\mathfrak{G}_{ic}^{a}$. The least decaying terms are in the third line of (2.40), which we can write as
$\begin{split}k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi}=&\widehat{k}_{c}{}^{a}\nabla_{i}\widehat{\Phi}+\widehat{k}_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}\widehat{k}_{i}{}^{j}\nabla_{b}\widehat{\Phi}\\ +(\widetilde{k}_{c}{}^{a}-H\delta_{c}{}^{a})\nabla_{i}\widehat{\Phi}&+(\widetilde{k}_{i}{}^{a}-H\delta_{i}{}^{a})\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}(\widetilde{k}_{i}{}^{j}-H\delta_{i}{}^{j})\nabla_{b}\widehat{\Phi}\\ &+\delta_{c}{}^{a}\nabla_{i}\widehat{\Phi}+\delta_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{ci}\nabla_{b}\widehat{\Phi}\end{split}$ (5.45)
Thus,
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}(k_{c}{}^{a}\nabla_{i}\widehat{\Phi}+k_{i}{}^{a}\nabla_{c}\widehat{\Phi}-g^{ab}g_{cj}k_{i}{}^{j}\nabla_{b}\widehat{\Phi})\|_{L^{2}(\Sigma_{s},g)}\\ \leq\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{-\frac{1}{2}Hs}e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}$ (5.46)
The rest of the terms in (2.40) satisfy better higher order estimates and are treated similarly.
For the inequality (5.43b), the least decaying terms come from the first line of (2.46):
$e^{Hs}e^{\ell Hs}\|\nabla^{(\ell)}(\widehat{\Phi}k_{l}{}^{l}k_{i}{}^{j}+\widetilde{\Phi}\widehat{\Phi}k_{i}{}^{j}-\Lambda\delta_{i}{}^{j}\widehat{\Phi})\|_{L^{2}(\Sigma_{s},g)}\leq Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}$
The rest of the terms in (2.46) satisfy better higher order estimates.
Finally, we turn to the estimate (5.43c) with $\mathfrak{F}$ given by (2.47). While for the first two terms in $\mathfrak{F}$,
$e^{3Hs}e^{2\ell Hs}\int_{\Sigma_{s}}f^{2}(t)|\nabla^{(\ell)}(\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}+\widehat{\Phi}\widetilde{k}_{i}{}^{j}\widehat{k}_{j}{}^{i})|_{g}^{2}e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq Ce^{Hs}\|\widehat{\Phi}\|^{2}_{W^{N-2,\infty}}e^{2Hs}\|\widehat{k}\|_{H^{N}}^{2}+C\|\widehat{k}\|^{2}_{W^{N-2,\infty}}e^{3Hs}\|\widehat{\Phi}\|_{H^{N}}^{2}\leq e^{-2Hs}\mathcal{E}_{N}(s)\,,$
we encounter the term decaying the least in $\Phi\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}=\widehat{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}+\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}$, and obtain:
$e^{3Hs}e^{2\ell Hs}\int_{\Sigma_{s}}f^{2}(t)|\nabla^{(\ell)}(\widetilde{\Phi}\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i})|_{g}^{2}e^{-3Hs}\mathrm{vol}_{g}\leq Ce^{Hs}\|\widehat{k}\|_{W^{N-2,\infty}}^{2}e^{2Hs}\|\widehat{k}\|_{H^{N}}^{2}\leq Ce^{-Hs}\mathcal{E}_{N}(s)\,.$
Here and above we have used Lemma 4.9 for the pointwise estimates and (4.12). The remaining terms in $\nabla^{(\ell)}\mathfrak{F}$ are estimated similarly now using the pointwise estimates of Lemma 3.3 for the reference metric.
Proposition 5.9.
Assume the bootstrap assumptions (4.12) are satisfied for some $N\geq 4$.
For all $s\in[s_{0},s_{b})$,
  1. (I)
    the error terms in (5.20a) and (5.20b) satisfy the estimates:
    $e^{\ell Hs}\|\mathrm{Error}_{\widehat{g},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}$ (5.47a)
    $e^{\ell Hs}\|\mathrm{Error}_{\widehat{g}^{-1},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}\,,$ (5.47b)
  2. (II)
    the error in (5.22) satisfies:
    $e^{\frac{3}{2}Hs}e^{\ell Hs}\|\mathrm{Error}_{\widehat{\Phi},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{\frac{3}{2}Hs}\|\widetilde{\mathfrak{I}}_{\Phi}\|_{W^{\ell,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\,.$ (5.48)
  3. (III)
    the error terms in (5.24), (5.26) satisfy:
    $e^{Hs}e^{\ell Hs}\|\mathrm{Error}_{\widehat{k},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{Hs}\|(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\|_{W^{\ell,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}$ (5.49)
    $e^{Hs}e^{\ell Hs}\|\mathrm{Error}_{\widehat{\Gamma},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq$ $\,Ce^{-\frac{1}{2}Hs}\sqrt{\mathcal{E}_{N}(s)}+Ce^{-\frac{1}{2}Hs}e^{\frac{3}{2}Hs}e^{\ell Hs}\|\nabla^{(\ell+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}$ (5.50)
  4. (IV)
    and the error terms in (5.28) satisfy:
    $e^{Hs}e^{\ell Hs}\|\mathrm{Error}_{\mathrm{div}\widehat{k},\ell}\|_{L^{2}(\Sigma_{s},g)}\leq\,Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}+C\|\widetilde{\mathfrak{C}}_{i}\|_{W^{\ell,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\,.$ (5.51)
Proof.
For (I) consider (5.20a),
$\mathrm{Error}_{\widehat{g},\ell}=$ $\nabla^{(\ell)}\mathrm{Error}_{\widehat{g}}+[\partial_{s},\nabla^{(\ell)}]\widehat{g}$ (5.52)
$(\mathrm{Error}_{\widehat{g}})_{ij}=$ $2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}$ (5.53)
For each term in $\mathrm{Error}_{\widehat{g}}$, we separate the differences, for example:
$\Phi g_{ja}\widehat{k}_{i}{}^{a}=\widehat{\Phi}\widehat{g}_{ja}\widehat{k}_{i}{}^{a}+\widetilde{\Phi}\widehat{g}_{ja}\widehat{k}_{i}{}^{a}+\widehat{\Phi}\widetilde{g}_{ja}\widehat{k}_{i}{}^{a}+\widetilde{\Phi}\widetilde{g}_{ja}\widehat{k}_{i}{}^{a}$ (5.54)
Then with $\ell$ derivatives falling on $\mathrm{Error}_{\widehat{g}}$, we estimate the highest order of derivatives of $\widehat{k}$, $\widehat{g}$, or $\widehat{\Phi}$ in energy, and apply the Sobolev inequalities of Lemma 4.9 to the lower orders, exploiting the decay of all quantities except $\widehat{g}$, in either norm:
$e^{\ell Hs}\|\nabla^{(\ell)}(\widehat{\Phi}\widehat{g}\widehat{k})\|_{L^{2}}\leq e^{-Hs}\Bigl{(}\|\widehat{\Phi}\|_{W^{N-2,\infty}}+\|\widehat{g}\|_{W^{N-2,\infty}}+\|\widehat{k}\|_{W^{N-2,\infty}}\Bigr{)}e^{Hs}\Bigl{(}\|\widehat{k}\|_{H^{N}}+\|\widehat{\Phi}\|_{H^{N}}\Bigr{)}\\ +e^{-Hs}\Bigl{(}e^{Hs}\|\widehat{\Phi}\|_{W^{N-2,\infty}}+e^{Hs}\|\widehat{k}\|_{W^{N-2,\infty}}\Bigr{)}\|\widehat{g}\|_{H^{N}}\leq Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}\,.$ (5.55)
For the terms involving derivatives of the reference metric, namely $\nabla^{(\ell)}\widetilde{\Phi}$ or $\nabla^{(\ell)}\widetilde{g}$, we also separate differences using $\nabla=\widehat{\nabla}+\widetilde{\nabla}$ thus introducing $\widehat{\Gamma}$ as a quantity that can be treated alongside $\widehat{k}$ as above. The remaining terms ~$\widetilde{\nabla}^{(\ell)}\widetilde{\Phi}$ and ~$\widetilde{\nabla}^{(\ell)}\widetilde{g}$ can always be estimated in $L^{\infty}$. For example,
$e^{\ell Hs}\|(\widetilde{\nabla}^{(\ell)}\widetilde{\Phi})\widehat{g}\widehat{k}\|_{L^{2}}\leq\|\widetilde{\Phi}\|_{W^{N,\infty}}\|\widehat{g}\|_{L^{\infty}}\|\widehat{k}\|_{L^{2}}\leq Ce^{-Hs}\sqrt{\mathcal{E}_{N}(s)}\,.$ (5.56)
In this way, all terms in $\nabla^{(\ell)}\mathrm{Error}_{\widehat{g}}$ can be as asserted. Together with Lemma 5.7, this implies (5.47).
For (II) consider (5.22). The commutator terms in $\mathrm{Error}_{\widehat{\Phi},\ell}$ are dealt with using Lemma 5.7, and the estimate for $\nabla^{(\ell)}\mathfrak{F}$ is given in Lemma 5.8. Also, the $L^{2}(\Sigma_{s},g)$ norm of $\nabla^{(\ell)}\widetilde{\mathfrak{J}}_{\Phi}$ can be replaced by the $W^{\ell,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ norm using Lemma 4.10 and the observation that $\widetilde{\mathfrak{J}}_{\Phi}$ is supported in $\{-1\leq t\leq 1\}$, cf. (2.48) and (3.20).
(III) now follows directly by employing Lemma 5.8, and the commutation estimates of Lemma 5.7, together with the bootstrap assumptions (4.12), and the Lemma 4.9.
(IV) follows similarly, using in addition Lemma 3.3 for the behavior of the reference variables in (5.28). ∎
5.4 Main energy estimates
In this section we derive the main energy estimates for the variables $\widehat{g},\widehat{g}^{-1},\widehat{k},\widehat{\Gamma}$, using the error estimates in Section 5.3 and the energy identities in Section 5.2.2
Proposition 5.10.
Assume that the bootstrap assumptions (4.12) are valid for some $N\geq 4$.
Then the following energy estimates hold for all $s\in[s_{0},s_{b})$:
  1. (I)
    For $\widehat{g},\widehat{g}^{-1}$,
    $\partial_{s}\bigl{(}\|\widehat{g}\|^{2}_{H^{N}(\Sigma_{s},g)}+\|\widehat{g}^{-1}\|^{2}_{H^{N}(\Sigma_{s},g)}\bigr{)}\leq Ce^{-Hs}\mathcal{E}_{N}(s),$ (5.57)
  2. (II)
    for $\widehat{\Phi}$,
    $\partial_{s}\big{\{}e^{3Hs}\|\widehat{\Phi}\|_{H^{N}(\Sigma_{s},g)}^{2}\big{\}}+e^{3Hs}\|\nabla\widehat{\Phi}\|_{H^{N}(\Sigma_{s},g)}^{2}\leq\\ \leq\,Ce^{-\frac{1}{2}Hs}\mathcal{E}_{N}(s)+Ce^{\frac{7}{2}Hs}\|\widetilde{\mathfrak{I}}_{\Phi}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2},$ (5.58)
  3. (III)
    and finally for $\widehat{\Gamma},\widehat{k}$,
    $\partial_{s}\big{\{}\frac{1}{3}e^{2Hs}\|\widehat{\Gamma}\|^{2}_{H^{N}(\Sigma_{s},g)}+e^{2Hs}\|\widehat{k}\|^{2}_{H^{N}(\Sigma_{s},g)}\big{\}}\leq\\ \leq\,Ce^{-\frac{1}{2}Hs}\mathcal{E}_{N}(s)+4e^{-\frac{1}{2}Hs}e^{3Hs}e^{2NHs}\|\nabla^{N+1}\widehat{\Phi}\|^{2}_{L^{2}(\Sigma_{s},g)}\\ +Ce^{\frac{5}{2}Hs}\|(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}+Ce^{\frac{1}{2}Hs}\|\widetilde{\mathfrak{C}}_{i}\|_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}^{2}\,.$ (5.59)
Proof.
For the derivation of the energy estimates we frequently use that the second fundamental form $k$ of the solution is diagonal up to a decaying remainder:
$|\Phi k_{i}{}^{j}-H\delta_{i}{}^{j}|\leq Ce^{-Hs}$ (5.60)
This is proven as follows. Since
$\Phi k-H\delta=\widehat{\Phi}\widehat{k}+\widetilde{\Phi}\widehat{k}+\widehat{\Phi}\widetilde{k}+\widetilde{\Phi}\widetilde{k}-H\delta$ (5.61)
and for the reference solution $\widetilde{\Phi}\widetilde{k}_{i}{}^{j}=H\delta_{i}{}^{j}+\mathcal{O}(e^{-2Hs})$ by Lemma 3.3, it remains to bound the differences. In view of the bootstrap assumptions, this follows from Lemma 4.9:
$e^{\frac{3}{2}Hs}|\widehat{\Phi}|+e^{Hs}|\widehat{k}|\leq C\varepsilon$ (5.62)
Each of the estimates (I)-(III) is derived by multiplying the corresponding energy identity in Section 5.2.2 with the weight $f^{2}(t)$ in (4.5) and then integrating on $\Sigma_{s}$ with respect to the volume form $e^{-3Hs}\mathrm{vol}_{g}$. A correction term is generated when $\partial_{s}$ falls on $e^{-3Hs}\mathrm{vol}_{g}$, namely
$\partial_{s}(e^{-3Hs}\mathrm{vol}_{g})=(\Phi\,\text{tr}_{g}k-3H)e^{-3Hs}\mathrm{vol}_{g},\qquad|\Phi\,\text{tr}_{g}k-3H|\leq Ce^{-Hs},$ (5.63)
which again follows from (5.60).
(I). Applying the above procedure to the energy identity of Lemma 5.29, we have
$\sum_{\ell\leq N}e^{2\ell Hs}\int_{\Sigma_{s}}\Bigl{\lvert}(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{g}\star\nabla^{(\ell)}\widehat{g}+(\Phi k-H\delta)\star\nabla^{(\ell)}\widehat{g}^{-1}\star\nabla^{(\ell)}\widehat{g}^{-1}\Bigr{\rvert}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq Ce^{-Hs}\mathcal{E}_{N}(s)$ (5.64)
and
$\sum_{\ell\leq N}e^{2\ell Hs}\int_{\Sigma_{s}}\Bigl{\lvert}\mathrm{Error}_{\widehat{g},\ell}\star\nabla^{(\ell)}\widehat{g}+\mathrm{Error}_{\widehat{g}^{-1},\ell}\star\nabla^{(\ell)}\widehat{g}^{-1}\Bigr{\}}\Bigr{\rvert}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq Ce^{-Hs}\mathcal{E}_{N}(s)$ (5.65)
which follow by Cauchy-Schwarz, the bound (5.60), and the error estimate (5.47) in Proposition 5.9.
(II). For the energy estimate (5.58), we repeat the previous argument, using instead the energy identity (5.34), and in addition integrate by parts the terms which take the form of a divergence:
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}\nabla^{i}\big{[}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big{]}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\\ =-\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{3Hs}e^{2\ell Hs}g^{ii^{\prime}}(\nabla_{i}\nabla^{(\ell)}\widehat{\Phi})\nabla^{(\ell)}\widehat{\Phi}\big{[}\nabla_{i^{\prime}}f^{2}(t)\big{]}e^{-3Hs}\mathrm{vol}_{g}\\ \leq\,Ce^{-Hs}\mathcal{E}_{N}(s)+\frac{1}{2}e^{-Hs}e^{3Hs}e^{2NHs}\|\nabla^{(N+1)}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}^{2}$ (5.66)
where we used that $|\partial f(t)|\leq(\alpha_{1}+\alpha_{2})f(t)$. The term with $(N+1)$ derivatives of $\widehat{\Phi}$ can be absorbed in the LHS thanks to the corresponding favorable term in (5.34). The stated estimate then follows from Proposition 5.9 (II) and Young’s inequality.
(III). For the last energy estimate (5.59), we argue similarly, using instead the energy identity (5.6). Integrating by parts the divergence terms in (5.36) produces error terms which are controlled as above:
$\sum_{\ell\leq N}\int_{\Sigma_{s}}(\mathrm{Div}_{\widehat{\Gamma},\widehat{k},\ell})f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq Ce^{-Hs}\mathcal{E}_{N}(s)+e^{-Hs}e^{3Hs}e^{2NHs}\|\nabla^{(N+1)}\widehat{\Phi}\|_{L^{2}}^{2}$ (5.67)
For the overall $e^{Hs}$ powers that appear in these estimates using Cauchy-Schwarz, it is useful to recall Lemma 4.7. For example,
$\int_{\Sigma_{s}}e^{2Hs}e^{2\ell Hs}\Phi g^{ii^{\prime}}g^{cc^{\prime}}g_{aa^{\prime}}\big{[}\nabla^{(\ell)}\widehat{\Gamma}_{i^{\prime}c^{\prime}}^{a^{\prime}}\star\nabla^{(\ell)}\widehat{k}_{c}{}^{a}\big{]}\nabla_{i}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq e^{-Hs}\|\widehat{\Gamma}\|_{H^{\ell}(\Sigma_{s},g)}\|\widehat{k}\|_{H^{\ell}(\Sigma_{s},g)}\,.$ (5.68)
Furthermore, we have in the second line of the RHS of (5.6),
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{2Hs}e^{2\ell Hs}\Bigl{[}\Phi g^{ab}g_{jj^{\prime}}\big{\{}\nabla^{j^{\prime}}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{j^{\prime}}\big{\}}\star\nabla^{(\ell)}\widehat{\Gamma}_{ab}^{j}\Bigr{]}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq e^{Hs}\Bigl{(}\|\nabla\widehat{\Phi}\|_{H^{N}(\Sigma_{s},g)}+\sum_{\ell\leq N}e^{\ell Hs}\|\mathrm{Error}_{\mathrm{div}\widehat{k},\ell}\|_{L^{2}(\Sigma_{s},g)}\Bigr{)}e^{Hs}\|\widehat{\Gamma}\|_{H^{N}(\Sigma_{s},g)}\\ \leq Ce^{-\frac{1}{2}Hs}\mathcal{E}_{N}(s)+e^{-\frac{1}{2}Hs}e^{3Hs}e^{2NHs}\|\nabla^{N+1}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}^{2}+Ce^{\frac{1}{2}Hs}\|\widetilde{\mathfrak{C}}_{i}\|^{2}_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\,.$ (5.69)
where we have used the error estimate of Proposition 5.9 (IV); similarly in the third line of RHS of (5.6)
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{2Hs}e^{2\ell Hs}\Bigl{[}\big{\{}\nabla^{c}\nabla^{(\ell)}\widehat{\Phi}+(\mathrm{Error}_{\mathrm{div}\widehat{k},\ell})^{c}\big{\}}\star\nabla^{(\ell)}\nabla_{c}\widehat{\Phi}\Bigr{]}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq e^{2Hs}\|\nabla\widehat{\Phi}\|_{H^{N}(\Sigma_{s},g)}^{2}+\sum_{\ell\leq N}e^{Hs}e^{\ell Hs}\|\mathrm{Error}_{\mathrm{div}\widehat{k},\ell}\|_{L^{2}(\Sigma_{s},g)}e^{Hs}\|\nabla\widehat{\Phi}\|_{H^{N}(\Sigma_{s},g)}\\ \leq 2e^{-Hs}e^{3Hs}e^{2NHs}\|\nabla^{N+1}\widehat{\Phi}\|_{L^{2}(\Sigma_{s},g)}^{2}+Ce^{-Hs}\mathcal{E}_{N}(s)+C\|\widetilde{\mathfrak{C}}_{i}\|^{2}_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\,.$ (5.70)
Finally, in the fourth line of (5.6) we can apply Proposition 5.9 (III) to obtain
$\sum_{\ell\leq N}\int_{\Sigma_{s}}e^{2Hs}e^{2\ell Hs}\Bigl{\{}\mathrm{Error}_{\widehat{\Gamma},\ell}\star\nabla^{(\ell)}\widehat{\Gamma}+\mathrm{Error}_{\widehat{k},\ell}\star\nabla^{(\ell)}\widehat{k}\Bigr{\}}f^{2}(t)e^{-3Hs}\mathrm{vol}_{g}\leq\\ \leq Ce^{-\frac{1}{2}Hs}\mathcal{E}_{N}(s)+e^{-\frac{1}{2}Hs}e^{3Hs}e^{2NHs}\|\nabla^{N+1}\widehat{\Phi}\|_{L^{2}(\Sigma_{S},g)}^{2}+Ce^{\frac{5}{2}Hs}\|(\widetilde{\mathfrak{J}}_{k})_{i}{}^{j}\|^{2}_{W^{N,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}$ (5.71)
This concludes the proof of the main energy estimates. ∎
6 Precise asymptotics of the perturbed solution
Now that we have established the global stability estimate (4.19), we can derive the precise asymptotic behavior of all variables.
Proposition 6.1.
The sharp estimate
$\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s},g)}+\|f(t)\widehat{\Phi}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-2Hs},$ (6.1)
holds for all $s\in[s_{0},+\infty)$. Moreover, the following expansions are valid for $\widehat{g}_{ij},\widehat{\Phi}$:
$\begin{split}\widehat{g}_{ij}(s,x)=\widehat{g}_{ij}^{\infty}(x)e^{2Hs}+\widehat{h}_{ij}(s,x),\\ \widehat{\Phi}(s,x)=\widehat{\Phi}^{\infty}(x)e^{-2Hs}+\widehat{\Psi}(s,x)\,,\end{split}$ (6.2)
where $\widehat{g}^{\infty}_{ij}(x),\widehat{h}_{ij}(s,x),\widehat{\Phi}^{\infty}(x),\widehat{\Psi}(s,x)$ satisfy:
$\begin{split}\|f(t)\widehat{g}^{\infty}(x)\|_{W^{N-4}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\qquad\|f(t)\widehat{h}(s,x)\|_{W^{N-4}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\\ \|f(t)\widehat{\Phi}^{\infty}(x)\|_{W^{N-6}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},\qquad\|f(t)\widehat{\Psi}(s,x)\|_{W^{N-6}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-4Hs}\,,\end{split}$ (6.3)
where the functions $\widehat{\Phi}^{\infty}(x),\widehat{\Psi}(s,x)$ are well-defined for $N\geq 6$.
Proof.
Using the global energy estimate (4.19) to control the RHS of (2.44), we deduce that
$\|f(t)(\partial_{s}\widehat{k}+3H\widehat{k})\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-\frac{3}{2}Hs}.$
Hence, we have
$\bigg{\|}\int_{s_{0}}^{s}\partial_{\tau}(f(t)e^{3H\tau}\widehat{k})\mathrm{d}\tau\bigg{\|}_{W^{N-4,\infty}(\Sigma_{s},g)}\leq$ $\,\int_{s_{0}}^{s}e^{3H\tau}\|f(t)(\partial_{\tau}\widehat{k}+3H\widehat{k})\|_{W^{N-4,\infty}(\Sigma_{\tau},g)}\mathrm{d}\tau\leq C\mathring{\varepsilon}e^{\frac{3}{2}Hs}$
$\Rightarrow\quad\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq$ $\,e^{-3Hs}e^{3Hs_{0}}\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s_{0}},g)}+C\mathring{\varepsilon}e^{-\frac{3}{2}Hs}\leq C\mathring{\varepsilon}e^{-\frac{3}{2}Hs},$
for all $s\in[s_{0},+\infty)$. Going back to the equation (2.45) for $\widehat{\Phi}$, we employ the latter improved estimate for $\widehat{k}$, together with the global estimate (4.19) to infer that
$\|f(t)(\partial_{s}\widehat{\Phi}+2H\widehat{\Phi})\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-3Hs}.$
Repeating the above argument, integrating in $[s_{0},s]$, gives
$\|f(t)\widehat{\Phi}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-2Hs}.$
Using the latter to bound the RHS of (2.44) once more, we obtain the improved estimate
$\|f(t)(\partial_{s}\widehat{k}+3H\widehat{k})\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-2Hs}.$
Integrating in $[s_{0},s]$ and repeating the above argument gives
$\|f(t)\widehat{k}\|_{W^{N-4,\infty}(\Sigma_{s},g)}\leq C\mathring{\varepsilon}e^{-2Hs},$
which completes the proof of (6.1).
Next, we employ the already derived (6.1), together with (4.19), to estimate the RHS of (2.37):
$\|f(t)(\partial_{s}\widehat{g}-2H\widehat{g})\|_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon},$
for all $(s,x)\in\mathcal{M}$. Hence, it follows that
$\bigg{\|}\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{-2Hs}\widehat{g})\mathrm{d}s\bigg{\|}_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\qquad s_{1}<s_{2}.$
This implies that $e^{-2Hs}\widehat{g}_{ij}$ has a $W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ limit, as $s\rightarrow+\infty$, denoted by $\widehat{g}_{ij}^{\infty}(x)$.
On the other hand, integrating in $[s,+\infty)$ gives:
$\bigg{\|}\int_{s}^{+\infty}\partial_{s}(f(t)e^{-2Hs}\widehat{g})ds\bigg{\|}_{W^{N-4,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs}$ (6.4)
Then with $\widehat{h}_{ij}=\widehat{g}_{ij}-\widehat{g}_{ij}^{\infty}e^{2Hs}$, it follows
[…]
The expansion for $\widehat{\Phi}$ is derived similarly, using the equation (2.45). The refined bounds (6.1), together with the global estimate, imply that
$\|f(t)(\partial_{s}\widehat{\Phi}+2H\widehat{\Phi})\|_{W^{N-6}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-4Hs}.$
Hence, it follows that
$\bigg{\|}\int_{s_{1}}^{s_{2}}\partial_{s}(f(t)e^{2Hs}\widehat{\Phi})\mathrm{d}s\bigg{\|}_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq$ $\,C\mathring{\varepsilon}e^{-2Hs_{1}},\qquad s_{1}<s_{2}.$
Hence, $e^{2Hs}\widehat{\Phi}$ has a $W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})$ limit, as $s\rightarrow+\infty$, denoted by $\widehat{\Phi}^{\infty}(x)$. Moreover, with $\widehat{\Psi}=\widehat{\Phi}-\widehat{\Phi}^{\infty}e^{-2Hs}$, integrating in $[s,+\infty)$ gives:
$\bigg{\|}\int_{s}^{+\infty}\partial_{s}(f(t)e^{2Hs}\widehat{\Phi})\mathrm{d}s$ $\bigg{\|}_{W^{N-6,\infty}(\mathbb{R}\times\mathbb{S}^{2},\mathring{g})}\leq C\mathring{\varepsilon}e^{-2Hs}$
[…]
This completes the proof of the proposition. ∎